### v1

# A Semantics-Aware Hierarchical Self-Supervised Approach to Classification of Remote Sensing Images

Abstract
Deep learning has become increasingly important in remote sensing image classification due to its ability to extract semantic information from complex data. Classification tasks often include predefined label hierarchies that represent the semantic relationships among classes. However, these hierarchies are frequently overlooked, and most approaches focus only on fine-grained classification schemes. In this paper, we present a novel Semantics-Aware Hierarchical Consensus (SAHC) method for learning hierarchical features and relationships by integrating hierarchy-specific classification heads within a deep network architecture, each specialized in different degrees of class granularity. The proposed approach employs trainable hierarchy matrices, which guide the network through the learning of the hierarchical structure in a self-supervised manner. Furthermore, we introduce a hierarchical consensus mechanism to ensure consistent probability distributions across different hierarchical levels. This mechanism acts as a weighted ensemble being able to effectively leverage the inherent structure of the hierarchical classification task. The proposed SAHC method is evaluated on three benchmark datasets with different degrees of hierarchical complexity on different tasks, using distinct backbone architectures to effectively emphasize its adaptability. Experimental results show both the effectiveness of the proposed approach in guiding network learning and the robustness of the hierarchical consensus for remote sensing image classification tasks.
I Introduction
Image classification and semantic segmentation are fundamental tasks in image processing. While Deep Learning (DL) methods have significantly advanced the field by learning powerful feature representations, standard approaches often treat classification as a flat, single-level task. However, many applications involve target categories having a natural hierarchical structure, exhibiting multiple levels of semantic granularity (e.g., object taxonomies, language, land-cover systems) [1, 2, 3]. Neglecting this inherent structure discards valuable semantic information and can lead to suboptimal or inconsistent predictions [4]. This is particularly relevant when dealing with label ambiguity or variations in image quality (such as resolution or artifacts) that may only permit reliable classification at coarser semantic levels.
A key challenge is integrating hierarchical information effectively into the learning process. Traditional methods frequently rely on post-processing steps or fixed rule-based systems to enforce hierarchical consistency, which limit their ability to learn complex, data-driven relationships between levels [5, 6, 7, 8]. Furthermore, constructing large-scale datasets with consistent multi-level annotations is often difficult [9, 10]. This requires methods capable of handling data labelled at various levels within the hierarchy. Although standardized hierarchical classification schemes exist in several domains, e.g., COoRdination of INformation on the Environment (CORINE) and Land Cover Classification System (LCCS) in Remote Sensing (RS), most current DL models do not explicitly leverage this structure during training.
Most hierarchical approaches focus on extracting features at multiple levels of granularity without explicitly modeling the hierarchical relationships intrinsic to the task [11, 12]. This poses a limitation, as the hierarchy arises from the problem domain itself rather than from specific architectural design choices. Accordingly, an effective hierarchical learning principle should be generalizable and scalable across different DL architectures. However, existing methods often produce non-modular architectures with rigid structures, thereby limiting the flexibility in the DL backbone selection.
In this paper, we propose a novel image processing methodology to modularly improve scene classification and semantic segmentation through semantic relationships across categories. Our approach leverages the hierarchical structure of standardized classification schemes by introducing a classification head for each hierarchical level. We further introduce trainable hierarchy matrices to model the relationships between these levels. To enforce consistency across the hierarchy, we define a multi-level consensus prediction that promotes agreement among the predictions at different levels. Additionally, we introduce a self-supervised consensus loss that implicitly learns the hierarchical relationships within the data. This approach allows the discovery of intricate sematic relationships that can deviate from the simple user-defined aggregation rules, leading to improved classification accuracy and semantic consistency. We demonstrate the effectiveness of our approach on different RS image classification tasks, including semantic segmentation of Satellite Image Time Series (SITS) of multispectral acquisitions and Very-High Resolution (VHR) RS scene classification tasks.
The key contributions of this work can be summarized as follows:
  1. 1.
    We define a hierarchical image processing approach to semantic segmentation and scene classification that integrates multi-level classification modules with DL backbones, along with trainable projection matrices to model hierarchical semantic relationships.
  2. 2.
    We propose a novel self-supervised consensus loss function that enforces consistency across different hierarchical levels, enhancing the semantic coherence of the classification results.
  3. 3.
    The proposed method is general and can be applied to any DL architecture and hierarchical classification scheme, offering flexibility in defining architecture details.
  4. 4.
    We demonstrate the effectiveness of the proposed approach on three RS datasets that include images with varying resolutions and different complexities in their hierarchical relationships.
The remainder of this paper is organized as follows. Section II reviews related works on hierarchical semantic segmentation. Section III details the proposed hierarchical consensus approach. The datasets and experimental setup are described in Section IV, while Section V presents and discusses the experimental results. Finally, Section VI concludes the paper and outlines future research directions.
II Related Works
Image classification is a fundamental task in computer vision and image processing, with DL models achieving state-of-the-art performance across various domains. Standard approaches typically treat classification as a “flat” problem in a traditional single-granularity classification, predicting probabilities over a set of mutually exclusive classes without explicitly considering potential semantic relationships or hierarchical structures among them. However, many real-world datasets possess inherent taxonomies where classes can be organized into hierarchies (e.g., species classification, object recognition with super-/sub-categories). Ignoring this structure can lead to suboptimal performance, as the model is not able to leverage prior knowledge about class relationships. Consequently, standard loss functions inherently treat all misclassifications equally. Without an explicit hierarchical representation, they fail to differentiate between errors involving closely related classes and those involving semantically distant ones.
Leveraging hierarchical label structures during model training and inference has emerged as a promising direction to improve accuracy and semantic understanding. Early work demonstrated that standard Convolutional Neural Networks implicitly capture some hierarchical information, with shallower layers grouping high-level categories and deeper layers specializing in finer distinctions [13]. Building on this, researchers have explored explicit methods to incorporate hierarchy. Typically, this is achieved by mapping the label hierarchy to network architectures or by introducing loss functions that handle the hierarchical correlation between different granularity tasks [14, 15, 16]. Bertinetto et al. [17] proposed a level-wise hierarchical cross-entropy loss based on conditional probabilities within the hierarchy and used mapping functions to embed class relationships. Similarly, Garg et al. [18] introduced the Hierarchy Aware Features (HAF), incorporating multiple losses, including minimizing the Jensen-Shannon Divergence (JSD) between aggregated predictions at different hierarchy levels and enforcing geometric consistency in the feature space, aiming to reduce the severity of misclassifications. In [19], the authors proposed hierarchy and exclusion graphs to model semantic relations between two labels assigned to the same objects. They considered exclusions, overlaps, and subsumptions, and provided a detailed theoretical analysis of the structure of the hierarchical problem. Building on this work, a combinatorial loss is proposed in [20], which addresses the scenario where samples might be labelled at different levels of granularity. The loss is designed to maximize the marginal probability of the observed ground truth label by aggregating information from related labels in the hierarchy. The authors introduced a Hierarchical Residual Network (HRN) architecture, where features from parent levels act as residual connections to inform child-level features, explicitly facilitating hierarchical feature interaction.
While these methods demonstrate the benefits of hierarchy-aware learning, they often rely on the assumption that a predefined and accurate hierarchy is available, where no hierarchy violation is possible. This assumption may not hold in all scenarios, as real-world taxonomies can be noisy, incomplete, or subjective. Furthermore, many existing techniques require explicit guidance or architectural modifications tailored to a specific fixed hierarchy [21], potentially limiting their flexibility and ability to adapt if the underlying hierarchical understanding evolves or is uncertain. For example, some approaches decompose the problem into sequential decisions or binary classifications guided by the known structure [22, 23, 21], rather than allowing the model to learn or refine hierarchical relationships implicitly from the data. The challenge remains to develop methods that can effectively exploit hierarchical semantics, even when the hierarchy is imperfect, labels are provided at varying granularity levels, or needs to be partially inferred in a robust and data-driven manner.
III Proposed Methodology
In this work, we address the above-mentioned challenges by proposing a novel deep learning framework for hierarchy-aware image classification. Our approach models the hierarchical semantic relationships within image data by integrating hierarchical loss components with a robust consensus mechanism. This allows the model to leverage potential hierarchical structures effectively while mitigating the impact of potential inaccuracies or inconsistencies in the predefined taxonomy. Specifically, our method aims to learn and exploit hierarchical relationships in a more implicit, self-supervised fashion within a hierarchy-aware architecture, reducing the reliance on rigid, externally defined structure.
III-A Hierarchical Problem Formulation
Let us denote column vectors by lowercase bold letters (e.g., $\mathbf{v}$) and matrices by uppercase bold letters (e.g., $\mathbf{B}$). The $i$th component of a vector $\mathbf{v}$ is denoted by $v_{i}$, while the $i$th row and $j$th column of a matrix $\mathbf{B}$ are denoted by $\mathbf{B}_{i,\cdot}$ and $\mathbf{B}_{\cdot,j}$, respectively. Then, let us consider a user-defined hierarchy label tree with $H$ levels, where the leaves represent the finest level of detail. We denote each hierarchical level as $h\in[1,2,\dots,H]$, ranging from the coarsest level $h=1$ to the finest one $h=H$. The root node $h=0$ is not considered, as it represents the trivial case in which all classes are mapped into a single class. Let $\mathbf{x}\in{\mathbb{R}^{b}}$ be the input vector, where $b$ is the number of input features. Its associated labels are $y^{h}\in\Omega^{h}$, where $\Omega^{h}=\{\omega^{h}_{1},\dots,\omega^{h}_{|\Omega^{h}|}\}$ is the set of labels at hierarchical level $h$, with $|\Omega^{h}|$ number of labels. For every $h_{1}<h_{2}$, we have that each fine-grained class $\omega_{j}^{h_{2}}$ is mapped to a coarse-grained class $\omega_{i}^{h_{1}}$, and $|\Omega^{h_{1}}|<|\Omega^{h_{2}}|$. We define this mapping relationship by $\omega_{j}^{h_{2}}\implies\omega_{i}^{h_{1}}$. Let $\mathcal{D}=\{(\mathbf{x}_{n},Y_{n})|n=1,\dots,N\}$ be the training set, where for each input feature vector $\mathbf{x}_{n}$, we have an associated set $Y_{n}=\{y_{n}^{1},\dots,y_{n}^{h},\dots,y_{n}^{H}\}$, which denotes the ground truth labels at all $H$ hierarchical levels, for a total number of $N$ training samples. The labels in the dataset satisfy a nested label structure, where each class at a finer level is contained within a class at a coarser level.
III-B Fully-Supervised Multi-Level Learning
The proposed methodology can be applied to any backbone network by simply introducing the required additional hierarchical classification modules. We refer to each of the hierarchical classification modules $\mathbf{g}^{h}:{\mathbb{R}^{d}\rightarrow\mathbb{R}^{|\Omega^{h}|}}$ on top of a given backbone network as the composition $\mathbf{g}^{h}\circ\mathbf{f}:{\mathbb{R}^{b}\rightarrow\mathbb{R}^{|\Omega^{h}|}}$, where $\mathbf{f}:{\mathbb{R}^{b}\rightarrow\mathbb{R}^{d}}$ is the backbone feature extractor. Fig. 1 provides a high-level overview of the considered multi-level classification approach.
Figure 1: Overview of the proposed hierarchical structure within the backbone (left) and associated hierarchical label tree (right). Different classification modules are trained to classify different hierarchical levels from the features extracted by the backbone.
Each classifier $\mathbf{g}^{h}$ estimates class-posterior logits. Let $\hat{\mathbf{p}}^{h}(\mathbf{x})\in\mathbb{R}^{|\Omega^{h}|}$ be the vector of estimated class-posterior probabilities at level $h$ given $\mathbf{x}$, i.e., $\hat{\mathbf{p}}^{h}(\mathbf{x})=[\hat{P}(\omega_{1}^{h}|\mathbf{x}),\dots,\hat{P}(\omega_{|\Omega^{h}|}^{h}|\mathbf{x})]^{\mathsf{T}}$, where $\mathsf{T}$ refers to the transpose operator. Then, the class-posterior probabilities $\hat{\mathbf{p}}^{h}$ for the input sample $\mathbf{x}$ at level $h$ can be retrieved as follows:
$\hat{\mathbf{p}}^{h}\left(\mathbf{x}\right)=\boldsymbol{\sigma}\left[\mathbf{g}^{h}\left(\mathbf{f}\left(\mathbf{x}\right)\right)\right],$ (1)
where $\boldsymbol{\sigma}:{\mathbb{R}^{d}\rightarrow[0,1]^{d}}$ is the softmax operator, with $\sigma_{i}[\cdot]$ computed as follows:
$\sigma_{i}\left[\mathbf{g}\right]=\frac{\exp{g_{i}}}{\sum_{j}\exp{g_{j}}}\in\left[0,1\right].$ (2)
The training of $\mathbf{g}^{h}$ is done by computing the categorical cross-entropy (CCE) loss as follows:
$\begin{split}\mathcal{L}_{CCE}^{h}&=-\frac{1}{N}\sum_{n=1}^{N}\sum_{i=1}^{|\Omega^{h}|}\mathbbm{1}\left[y^{h}_{n}=\omega^{h}_{i}\right]\log\hat{P}^{h}\left(\omega^{h}_{i}|\mathbf{x}_{n}\right)\\ &=-\frac{1}{N}\sum_{n=1}^{N}\sum_{i=1}^{|\Omega^{h}|}\mathbbm{1}\left[y^{h}_{n}=\omega^{h}_{i}\right]\log\sigma_{i}\left[\mathbf{g}^{h}\left(\mathbf{f}\left(\mathbf{x}_{n}\right)\right)\right],\end{split}$ (3)
where $\mathbbm{1}\left[\cdot\right]$ serves as an indicator function and takes a value of one when the argument is true, else zero. We apply the CCE loss function to all hierarchical levels, obtaining the total CCE loss $\mathcal{L}_{CCE}$ as the sum of the level-wise losses:
$\mathcal{L}_{CCE}=\sum_{h=1}^{H}\lambda^{h}\mathcal{L}_{CCE}^{h}.$ (4)
As it has also been noted in previous works [24], adopting several CCE loss functions for diverse classification levels may negatively affect performance at the most granular classification level. To mitigate this behavior, we introduced the weighting terms $\lambda=\{\lambda^{1},\dots,\lambda^{H}\}$ to limit the impact of coarser levels, while still accounting for the additional losses.
III-C Hierarchy Matrices
The CCE loss functions at each hierarchical level lead to independent classifications, which are not constrained to follow the hierarchy definition formulated in the problem statement. To impose the hierarchical constraints, additional consensus predictions and self-supervised penalty terms are introduced. These penalty terms are defined through hierarchy projection matrices designed to map the correct relationships between the different layers.
Let us consider for simplicity the case of $H=2$ levels, with $h_{1}$ representing the coarse-grained labels and $h_{2}$ the fine-grained labels, i.e., $h_{1}<h_{2}$. We refer to $\Omega^{h_{1}}$ as the set of the macroclasses (i.e., the coarse-grained classes) and to $\Omega^{h_{2}}$ as the set of the microclasses (i.e., the fine-grained classes). We can define the hierarchy indicator matrix $\mathbf{I}^{h_{2},h_{1}}\in\{0,1\}^{|\Omega^{h_{2}}|\times|\Omega^{h_{1}}|}$, which represents the hierarchical relationship between level $h_{2}$ and $h_{1}$:
$\mathbf{I}^{h_{2},h_{1}}=\begin{bmatrix}I_{1,1}&I_{1,2}&\dots&I_{1,|\Omega^{1}|}\\ I_{2,1}&I_{2,2}&\dots&I_{2,|\Omega^{1}|}\\ \vdots&\vdots&\ddots&\vdots\\ I_{|\Omega^{2}|,1}&I_{|\Omega^{2}|,2}&\dots&I_{|\Omega^{2}|,|\Omega^{1}|}\end{bmatrix}$. (5)
This approach can be easily extended to a multi-level scenario, considering any given number $H$ of hierarchical levels. In this case, the total number of hierarchy matrices that can be determined is $((H-1)\cdot H)/2$, as also mappings of distant relationship levels can be defined (e.g., from the root of the label tree to the leaves, skipping intermediate levels) if needed. An example is illustrated in Fig. 2.
Figure 2: Illustration of both direct and distant hierarchical mappings, from fine-grained classes to coarse-grained classes.
The hierarchy indicator matrices are initialized with ones and zeros representing the user-defined mapping. Consequently, $I^{h_{2},h_{1}}_{i,j}$ assumes a value of one when the microclass $\omega_{i}^{h_{2}}$ is contained in the macroclass $\omega_{j}^{h_{1}}$ through the relationship $\omega_{i}^{h_{2}}\implies\omega_{j}^{h_{1}}$}, thus defining the matrix elements as follows:
$I^{h_{2},h_{1}}_{i,j}=\begin{cases}1&\text{if }{\omega_{i}^{h_{2}}\implies\omega_{j}^{h_{1}}}\\ 0&\text{otherwise}\end{cases}.$ (6)
It is important to note that, in the case of hard-hierarchies, each microclass corresponds to exactly one macroclass. However, this may not hold true as there might exist semantic overlaps in the hierarchy structure. Additionally, a distinction between user-defined aggregations and data-driven aggregations has to be made. While the first is defined a-priori for the problem at hand by the user, the latter is not known and must be inferred directly from the data, reflecting the true underlying structure of the problem considered. To take this into account, we can soften the hierarchical relationships by allowing the matrices to take any value between 0 and $1$. In this way, one microclass may be associated with multiple macroclasses with different confidences reflecting the semantic overlaps. To this purpose, we can define hierarchy joint distribution matrices $\mathbf{J}^{h_{2},h_{1}}\in{[0,1]}^{|\Omega^{h_{2}}|\times|\Omega^{h_{1}}|}$. Each term of the matrix $J^{h_{2},h_{1}}_{i,j}=P(\omega^{h_{1}}_{j},\omega^{h_{2}}_{i})$ models the probability of observing a relationship between microclass $\omega^{h_{2}}$ and macroclass $\omega^{h_{1}}$. This constraint relaxation allows us to better combine the user-defined aggregation problem with a data-driven centered architecture, by learning these matrices directly from the data.
III-D Self-Supervised Hierarchical Consensus Learning
Let us return to the general case of $H$ hierarchical levels, considering projections from source levels $h$ to a target hierarchical level $\tilde{h}$. Each hierarchical level acts independently of the others. To ensure that the network is hierarchy-aware, an additional constraint based on the prediction associated to the target hierarchical level and the projected predictions from the other levels is defined. In particular, we take inspiration from the disagreement-based active learning field [25] to define the disagreement between the predictions calculated at each hierarchical level. For each level $\tilde{h}$, we consider a committee composed of all $H$ levels. All predictions and their projected versions can be viewed as “soft” votes, which define each committee member’s confidence. Then, we introduce a self-supervised loss that penalizes the committee members’ disagreement with the overall committee consensus, thus guiding the learning of data-driven relationships in $\mathbf{J}^{h,\tilde{h}}$ and providing further supervision to the network based on the predictions at the different levels.
III-D1 Hierarchy-based Projected Predictions
In order to generate the committee at each hierarchical level, the predictions first need to be projected to the other levels. For simplicity, let us consider the case of projection from source-level classes to target-level classes. Here we refer to estimated matrices as $\mathbf{\hat{J}}^{h,\tilde{h}}\in{[0,1]}^{|\Omega^{h}|\times|\Omega^{\tilde{h}}|}$. The projected probability distribution $\hat{P}^{h\rightarrow\tilde{h}}{(\omega^{\tilde{h}}_{j}|\mathbf{x})}$ from source-level classes to target-level classes can be obtained from the probability distribution over the source-level classes $\hat{P}^{h}{(\omega^{h}_{i}|\mathbf{x})}$ and the hierarchy projection probabilities $\hat{P}{(\omega^{\tilde{h}}_{j}|\omega^{h}_{i})}$ through the following relationship:
$\hat{P}^{h\rightarrow\tilde{h}}{\left(\omega^{\tilde{h}}_{j}|\mathbf{x}\right)}=\sum_{i=1}^{|\Omega^{h}|}\hat{P}{\left(\omega^{\tilde{h}}_{j}|\omega^{h}_{i}\right)}\cdot\hat{P}^{h}{\left(\omega^{h}_{i}|\mathbf{x}\right)},$ (7)
where the projection probabilities $\hat{P}{(\omega^{\tilde{h}}_{j}|\omega^{h}_{i})}$ are directly derived by row-normalizing the estimated joint matrices $\mathbf{\hat{J}}^{h,\tilde{h}}$, i.e., we can define hierarchy projection matrices $\mathbf{\hat{M}}^{h,\tilde{h}}\in{[0,1]}^{|\Omega^{h}|\times|\Omega^{\tilde{h}}|}$ such that $\hat{M}^{h,\tilde{h}}_{i,j}=\hat{P}(\omega^{\tilde{h}}_{j}|\omega^{h}_{i})$. In general vector form, we can write the class-posterior probability vector $\hat{\mathbf{p}}^{h\rightarrow\tilde{h}}{(\mathbf{x})}$ projected from level $h$ to level $\tilde{h}$ as follows:
$\hat{\mathbf{p}}^{h\rightarrow\tilde{h}}{(\mathbf{x})}=(\hat{\mathbf{M}}^{h,\tilde{h}})^{\mathsf{T}}\cdot\hat{\mathbf{p}}^{h}{\left(\mathbf{x}\right)}.$ (8)
To ensure numerical stability and maintain better control over gradient magnitudes, we consider the computation of all the losses directly in the log domain. First, we compute the class-posterior log probabilities of classification modules $\mathbf{g}^{h}$ from Eq. 2:
$\log{\hat{P}^{h}}{\left(\omega_{i}^{h}|\mathbf{x}\right)}=g_{i}^{h}-\text{LSE}{\left(\mathbf{g}^{h}\right)},$ (9)
where $\text{LSE}:{\mathbb{R}^{|\Omega^{h}|}\rightarrow\mathbb{R}}$ is the log-sum-exp (LSE) operator:
$\text{LSE}\left(\mathbf{g}^{h}\right)=\log\sum_{i=1}^{|\Omega^{h}|}{\exp{\left(g_{i}^{h}\right)}}.$ (10)
Then, we can also define the projected predictions $\hat{P}^{h\rightarrow\tilde{h}}{(\omega_{j}^{\tilde{h}}|\mathbf{x})}$ from level $h$ to level $\tilde{h}$ in the log domain, starting from the definition in the natural domain. Applying the logarithm to both sides of Eq. 7, we obtain the following:
$\log{\left(\hat{P}^{h\rightarrow\tilde{h}}{\left(\omega^{\tilde{h}}_{j}|\mathbf{x}\right)}\right)}=$
$\hskip-70.0001pt=\log{\left(\sum_{i}\hat{M}^{h,\tilde{h}}_{i,j}\cdot\hat{P}^{h}{\left(\omega^{h}_{i}|\mathbf{x}\right)}\right)}$
$\hskip-70.0001pt=\log{\left(\sum_{i}\hat{M}^{h,\tilde{h}}_{i,j}\cdot\frac{\exp{g_{i}^{h}}}{\sum_{k}\exp{g_{k}^{h}}}\right)}$
$\hskip-70.0001pt=\log{\left(\sum_{i}\exp\left({g_{i}^{h}+\log{\hat{M}^{h,\tilde{h}}_{i,j}}}\right)\right)}-\log{\sum_{k}\exp{g_{k}^{h}}}.$ (11)
This can finally be rewritten as the combination of two LSE operations, yielding the projected log-probability distribution at level $\tilde{h}$:
$\log{\left(P^{h\rightarrow\tilde{h}}\left(\omega_{j}^{\tilde{h}}|\mathbf{x}\right)\right)}=\\ =\text{LSE}\left(\mathbf{g}^{h}+\log{\left(\mathbf{\hat{M}}^{h,\tilde{h}}_{\cdot,j}\right)}\right)-\text{LSE}\left(\mathbf{g}^{h}\right).$ (12)
To further simplify, hierarchy projection matrices can be defined directly in the log domain, i.e., we learn log joint probabilities $\mathbf{L}^{h,\tilde{h}}=\log{\hat{\mathbf{J}}^{h,\tilde{h}}}\in\mathbb{R}^{|\Omega^{h}|\times|\Omega^{\tilde{h}}|}$. Then, we compute $\log{\hat{\mathbf{M}}^{h,\tilde{h}}}$ as follows:
[…]
which is equivalent to the row-normalization of $\hat{\mathbf{J}}^{h,\tilde{h}}$ in the log domain. The projected logits $\log{\hat{\mathbf{p}}}^{h\rightarrow\tilde{h}}$ can be used to calculate additional CCE losses that can contribute to the final loss [26]. However, this does not explicitly model the hierarchical consensus at the different hierarchical levels, requiring an additional step to ensure the joint classification of the classes.
III-D2 Semantics-Aware Hierarchical Consensus
Given the projected prediction at each hierarchical level, we can now define a Hierarchical Consensus (HC) class probability vector $\hat{\mathbf{p}}^{\tilde{h}}_{HC}{(\mathbf{x})}$:
$\hat{\mathbf{p}}^{\tilde{h}}_{HC}{\left(\mathbf{x}\right)}=$
$\hskip-50.00008pt=\frac{1}{Z^{\tilde{h}}}\sqrt[H]{\hat{\mathbf{p}}^{\tilde{h}}{\left(\mathbf{x}\right)}\prod_{h\neq\tilde{h}}\hat{\mathbf{p}}^{h\rightarrow\tilde{h}}{\left(\mathbf{x}\right)}}$
$\hskip-50.00008pt=\frac{1}{Z^{\tilde{h}}}\sqrt[H]{\hat{\mathbf{p}}^{\tilde{h}}{\left(\mathbf{x}\right)}\prod_{h\neq\tilde{h}}(\mathbf{\hat{M}}^{h,\tilde{h}})^{\mathsf{T}}\cdot\hat{\mathbf{p}}^{h}{\left(\mathbf{x}\right)}},$ (14)
where $\hat{\mathbf{p}}^{\tilde{h}}_{HC}$ is the HC class-posterior probability defined as the geometric mean of the predictions at the target hierarchical level $\tilde{h}$ and the projected predictions of all the other $h$ levels, and $Z^{\tilde{h}}$ is a normalization constant to ensure that $\mathbf{p}^{\tilde{h}}_{HC}$ sums to one. This is equivalent to computing the average in the log domain [27]:
$g^{\tilde{h}}_{HC_{j}}$ $=\frac{1}{H}\left[\log{\hat{P}^{\tilde{h}}{\left(\omega^{\tilde{h}}_{j}|\mathbf{x}\right)}}+\sum_{h\neq\tilde{h}}\log{\hat{P}^{h\rightarrow\tilde{h}}{\left(\omega^{\tilde{h}}_{j}|\mathbf{x}\right)}}\right]$
[…]
$+\sum_{h\neq\tilde{h}}\left[\text{LSE}\left(\mathbf{g}^{h}+\log{\mathbf{\hat{M}}_{\cdot,j}^{h,\tilde{h}}}\right)-\text{LSE}\left(\mathbf{g}^{h}\right)\right]\bigg\},$
$\hskip-50.00008pt\log{\hat{\mathbf{p}}^{\tilde{h}}_{HC}{\left(\mathbf{x}\right)}}=\mathbf{g}^{\tilde{h}}_{HC}-\text{LSE}{\left(\mathbf{g}^{\tilde{h}}_{HC}\right)}$, (15)
where $\text{LSE}{(\mathbf{g}^{\tilde{h}}_{HC})}=\log{Z^{\tilde{h}}}$. While not directly supervised, HC can serve as an in-place replacement for the original classification modules during inference, providing a fully semantics-aware prediction in the downstream task.
III-D3 Self-Supervised Consensus Loss
After calculating the consensus, both the prediction at the target level and the projected predictions are compared with the HC, defining an additional loss based on the divergence from the HC. To this purpose, we employ the JSD, which is defined as follows:
$\text{JSD}^{h,\tilde{h}}\left(\hat{\mathbf{p}}^{\tilde{h}}_{HC}||\hat{\mathbf{p}}^{h\rightarrow\tilde{h}}\right)=\\ =\frac{1}{2}\left(\text{KLD}\left(\hat{\mathbf{p}}^{\tilde{h}}_{HC}||\mathbf{q}\right)+\text{KLD}\left(\hat{\mathbf{p}}^{h\rightarrow\tilde{h}}||\mathbf{q}\right)\right),$ (16)
[…]
$\mathcal{L}_{HC}=\frac{1}{N}\sum_{n=1}^{N}\sum_{\tilde{h}=1}^{H}\frac{1}{\log{|\Omega^{\tilde{h}}|}}\sum_{h=1}^{H}{\text{JSD}^{h,\tilde{h}}\left(\mathbf{x}_{n}\right)},$ (17)
[…]
III-E Overall Loss Function
Lastly, the losses are averaged and summed to $\mathcal{L}_{CCE}$ (Eq. 4):
$\mathcal{L}=\mathcal{L}_{CCE}+\lambda_{HC}\mathcal{L}_{HC},$ (18)
where $\lambda_{HC}$ is a metaparameters tuning the contribution of the HC loss. It is important to note that $\mathcal{L}_{HC}$ does not require supervision, as it is simply a loss representing the divergence from the HC. Though the HC prediction is used within the penalty loss during the training phase, we consider the HC prediction at each hierarchical level as the final predictions for inference. This allows the architecture to leverage the semantic relationships learnt during the self-supervised training phase and to perform hierarchically consistent predictions.
IV Dataset Description and Experiment Setup
This Section provides a description of the datasets employed in this study. Then, the experimental setup, including the preprocessing phase and the benchmark methodologies adopted for evaluating the DL models, are described.
IV-A NWPU-RESISC45
To assess the performance of the proposed methodology on VHR images, where spatial features play a prominent role, we considered the NWPU-RESISC45 [28] benchmark as a first dataset. NWPU-RESISC45 is composed of $31.500$ RGB multi-resolution images for scene classification of VHR images categorized into 45 different classes at the finest-level of detail, and exhibits a high within-class diversity and between-class similarity. The dataset is completely balanced, i.e., each class has the same number of samples at the fine-grained level. We defined a custom-hierarchy based on the subdivision defined in MillionAID [29], resulting in 8, 24, and 45 classes, corresponding to coarse-grained, intermediate, and fine-grained classes, respectively.
IV-B ELU Dataset
Figure 3: A semicircular hierarchical label tree representation of the Emilia LU dataset at the fine-grained level of the hierarchy. The leaves corresponding to the fine-grained classes are shown in green, while the nodes for intermediate classes are depicted in blue, and those for coarse-grained classes are shown in red. The root node is represented in black.
To further validate the performance of the proposed methodology, we additionally considered two multitemporal multispectral datasets differing in terms of complexity in the hierarchical structure. The first dataset [^1] Land Use of Emilia (2020). Accessed on May $13^{\text{th}}$ $2025$. consists of a publicly available LU map of $2020$ over the Emilia Romagna region, Italy. The maps were defined considering the standard CORINE Land Cover (CLC) aggregation at three levels, with the third one further divided, resulting in an additional fourth level with an increased semantic detail. For the purpose of this analysis, we evaluated the hierarchical approach on three levels of hierarchy, consisting in the first two levels of the CLC and the additional finest level of detail, here referred to as coarse-grained, intermediate and fine-grained, respectively. The dataset has been built considering orthophotos and infrared images, and shows a minimum mapping unit of $0.16$ hectares, with a minimum of $7\text{m}$ meters for linear structures. Although it has a good positional accuracy of $5\text{m}$, the data may exhibit errors. Additionally, it represents LU classes, where the same Land-Cover (LC) class could be associated with different semantic classes, such as the fine-grained classes “Commercial airports”, “Sport flight airports and heliports”, and “Military airports”. We considered the Emilia Land Use (ELU) dataset corresponding to the Military Grid Reference System (MGRS) Sentinel-2 (S2) tile 32TPQ ($10980\times 10980$ pixels, for a total area of $110\times 110$ $km^{2}$). In the analyzed tile, the three hierarchical levels show $5$, $18$, and $80$ classes at the coarse-grained, intermediate, and fine-grained levels, respectively. Please note that such an elevated number of fine-grained classes is not commonly approached in Earth Observation (EO) classification problems, where the majority of studies focus on the intermediate or coarse-grained classification task. Nonetheless, this level of precision allows a better monitoring of the class of interest and the extraction of information-rich content. Due to the high number of classes and, consequently, the large size of the hierarchical tree, we refer to this hierarchy as a complex hierarchy. A semicircular tree representing the hierarchical structure of the dataset is shown in Fig. 3.
IV-C ESA HRLC CCI Dataset
The second multitemporal multispectral dataset [^2] ESA High-Resolution Land Cover Climate Change Initiative Project CEDA Archive site. Accessed on May $13^{\text{th}}$ $2025$. has been derived under the High Resolution Land Cover (HRLC)- Climate Change Initiative (CCI) project by European Space Agency (ESA) [30]. The dataset includes LC classes organized according to the LCCS nomenclature, following the classification at the finest level of details. In this case, the hierarchical structure is less rigid compared to the Emilia dataset, as the classes are less detailed and can almost be considered independent. In this analysis, we use the HRLC1010 maps, i.e., the $10\text{m}$ resolution maps for the year $2019$. The maps have been derived from machine learning approaches considering a joint approach based on multispectral (MS) S2 and Synthetic Aperture Radar (SAR) Sentinel-1 (S1) images [31]. The HRLC1010 maps have been validated by intercomparison with other available maps and with photointerpreted samples. For our analysis, we considered the dataset corresponding to the MGRS S2 tile 22KGV ($10980\times 10980$ pixels, for a total area of $110\times 110$ $km^{2}$) in Brazil, mostly due to the heterogeneous presence of different classes. Since the hierarchical structure is simpler compared to the Emilia LU dataset, we defined three hierarchical levels ad-hoc, based on the similarity of the classes at the finest level of detail. This led to the definition of $3$, 77, and $12$ classes for the different levels, corresponding to coarse-grained, intermediate, and fine-grained classes, respectively. The coarse-grained classes have been divided in “Woody vegetated areas”, “Herbaceous vegetated areas”, and “Non-vegetated areas”. The fine-grained classes have been aggregated at the intermediate level considering the seasonality (e.g., “Tree cover evergreen broadleaf” and “Tree cover deciduous broadleaf” are aggregated at intermediate level, while “Tree cover evergreen needleleaf” is still separated and aggregated at the coarse-grained level instead). Since the number of classes only shows a slight increment in the different layers, we refer to this hierarchy as a simple hierarchy.
IV-D S2 Preprocessing and Dataset Preparation
IV-D1 VHR dataset
The training, validation, and test splits were selected accordingly to [32], considering a split-ratio of $0.6$, $0.2$, and $0.2$, respectively. During training, we applied data augmentation techniques consisting in resize, random crops and random rotations of the images.
IV-D2 Multitemporal MS datasets
The input data consists of S2 MS satellite images provided in the Copernicus EO program. The S2 acquisitions considered have been preprocessed at Level-2A (L2A) and re-projected at $10\text{m}$/pixel resolution through a nearest neighbor approach. In the experiments, all the bands except the $60\text{m}$ resolution band were used. Due to the high-complexity of both classification problems, we collected two time series (TS) of S2 images, fully exploiting the temporal signature and allowing the detection of seasonal targets. The S2 TS considered for the Emilia Dataset consists of acquisitions ranging from January $1^{\text{st}}$ $2020$ to December $31^{\text{st}}$ $2020$, while the S2 TS for the HRLC-CCI dataset consists of TS ranging from January $1^{\text{st}}$ $2019$ to December $31^{\text{st}}$ $2019$. Both the S2 TSs have been processed to achieve more regular TSs, resulting in $12$ median monthly composites. This was achieved by computing the pixel-wise median for each band for each month. Before applying the median operator, we masked the samples associated with saturate or defective pixels, clouds, and cloud shadows. This approach allowed us to have regular, gap-free data in input to the DL models.
To extract a balanced training set, we considered an approach based on the heterogeneity of the input training patches. In particular, for both datasets, we divided the initial S2 monthly composite images into non-overlapping patches of $9\times 9$ pixels. After the initial subdivision, we collected the patches and calculated the mode, extracting the dominant class of each patch analyzed. Then, the patches have been randomly subdivided into training, validation, and test sets, by considering a minimum number of $N$ representative patches for each class in the analyzed classification scheme, with $N$ fixed to $1000$ for both the ELU and HRLC-CCI datasets. In the case of a minoritarian class had fewer patches than $N$, $50\%$ of its total were assigned to the training set, $25\%$ to the validation set, and $25\%$ to the test set. In this way, we collected a small training dataset representative of the entirety of the classification scheme, corresponding to $\sim 1.35\%$ ($1.631.616$ pixels) and to $\sim 0.22\%$ ($268.800$ pixels) of the total number of patches in the tile for the Emilia Dataset and the HRLC-CCI dataset, respectively. The validation set size has been determined by considering $25\%$ of the number of samples allocated for each class for the training set, while the remainder of the samples (i.e., the entirety of the patches in the tile not belonging to the training and validation sets) have been used to validate the proposed approach in the test set.
IV-E Design of Experiments
Instead of validating the proposed hierarchical approach on a single backbone, we demonstrated its broad applicability by employing diverse backbone architectures representing different paradigms in a robust manner. All tests were conducted five times, with different random initializations of the weights of the architecture. For reproducibility, predefined seed values were used for the random number generator (RNG). The reported results include the values of Overall Accuracy (OA) for both tasks, while mean F1 Score (mF1) was also employed in the SITS MS classification task, along with the standard deviation (calculated as an unbiased estimator). For each architecture, we considered the following approaches: (i) the baseline with only the fine-grained classification through standard CCE minimization, (ii) the HAF method [18], (iii) the HRN module [20] and (iv) the proposed Semantics-Aware Hierarchical Consensus (SAHC). We removed the Granularity-Specific Block (GSB) block (two parameter-heavy convolutions and fully connected layers) from the HRN approach, maintaining only the hierarchical module. This is to ensure a fair comparison of architectures with the same number of parameters and focus solely on the hierarchical representation of the problem.
The different strategies are evaluated on the VHR scene classification task and the multitemporal MS semantic segmentation tasks. For the VHR task, we considered three different backbone architectures focusing on three different paradigms: (i) EfficientformerV2 S2 [33] for transformer-based architectures, (ii) MobileNetV2 [34] for CNN-based architectures, and (iii) LWGANet [35] for hybrid architectures. For the MS SITS classification task, we considered three different backbone architectures designed to process temporal information, also including hybrid architectures to further examine the effectiveness of the methodologies in utilizing contextual information: (i) the encoder part of a Transformer [36], (ii) an adaptation to the RS scenario of a Swin Transformer originally intended for super-resolution of videos [37], and (iii) a Convolutional Long Short-Term Memory (ConvLSTM) [38]. For a detailed description of the layer topology, please refer to the cited papers.
For each backbone, we replaced the classification head with three hierarchical classification heads, each characterized by a different level of granularity. We adopted the AdamW optimizer for all the architectures and configurations. The metaparameters of each different configuration have been validated considering a Tree-structured Parzen Estimator (TPE) algorithm to efficiently explore the search space. The learning rate and the weight decay were sampled from log-uniform distributions $U_{\log}([10^{-6},10^{-3}])$. Considering the proposed SAHC approach, the weight $\lambda_{HC}$ associated to the HC loss has been sampled from a log-uniform distribution $U_{\log}([10^{-2},10^{2}])$, while the weights $\lambda^{h}$ have been set empirically to $0.3$, $0.2$, and $0.5$, to prioritize the fine-grained classification and reduce the influence of the intermediate and coarse-grained levels. To adjust the learning rate during training, we considered a cosine decay schedule. The weights associated to each model in the best performing epoch in terms of validation OA% and mF1% were used to conduct the inference on the test set, whose performance metrics are reported in the following section.
V Experimental Results
The experiments focus on validating the proposed SAHC approach over different backbones tasks, and datasets, comparing it to different standard and hierarchical approaches. We first focus on the results on the specific tasks, i.e., scene classification on the VHR dataset and semantic segmentation of MS SITS. Then, we demonstrate the effectiveness of each component of the SAHC approach in an ablation study considering the ELU dataset. Finally, we analyze the trained hierarchy projection matrices to identify the key hierarchical relationships learnt by the proposed approach.
V-A Case I: VHR Scene Classification
TABLE I: Comparison of the Overall Accuracy obtained by different hierarchical methodologies with different backbone architectures on the NWPU-RESISC45 dataset. The best results are reported in bold, second best are underlined. The related standard deviation is reported in brackets.
NWPU-RESISC45
Backbone Methodology OA (%)
LWGANet L2 [35] Baseline 92.50 (0.31)
HAF [18] 93.05 (0.29)
HRN [20] 93.01 (0.11)
SAHC (proposed) 94.69 (0.18)
MobileNet V2 2.5× [34] Baseline 93.70 (0.30)
HAF 93.67 (0.28)
HRN 93.58 (0.16)
SAHC (proposed) 94.19 (0.33)
EfficientformerV2 S2 [33] Baseline 90.41 (0.38)
HAF 90.76 (0.46)
HRN 91.16 (0.44)
SAHC (proposed) 91.72 (0.14)
Table I shows the OAs of the LWGANet, MobileNet, and EfficientFormer on the NWPU-RESISC45 dataset. The proposed SAHC achieves the best results, showing a higher OA% with all the different backbones. The HAF and HRN hierarchical modules show improvements only in the case of the transformer-based backbone (i.e., EfficientFormer) and the hybrid backbone (i.e., LWGANet), while the baseline approach without hierarchical modules performs better with the CNN-based approach (i.e., MobileNet). In contrast, the proposed SAHC shows an increment in OA% of $\sim 1.2\%$ compared to the baseline approach with both the transformer-based and the CNN-based architectures, while exhibits a greater improvement when considering the hybrid backbone, with an additional $\sim 2.2\%$ increase in mF1%. These findings highlight the contribution of the hierarchical module. Indeed, its application to the hybrid architecture results in higher performance compared to its application to the CNN model. This result is particularly interesting given that the baseline CNN model outperforms the LWGANet baseline. It is also worth noting that, in most cases, the SAHC shows a smaller variability compared to its competitors, ensuring stabler results across different random initializations of the architecture weights.
V-B Case II: MS SITS Semantic Segmentation
TABLE II: Comparisons of the accuracies obtained by different hierarchical methodologies with different backbone architectures on the ELU and HRLC-CCI datasets. The best results are reported in bold, second best are underlined. The related standard deviation is reported in brackets.
ELU Dataset HRLC-CCI Dataset
Backbone
Methodology
mF1 (%)
OA (%)
mF1 (%)
OA (%)
Swin Transformer [37] Baseline (no hierarchy) 33.42 (0.44) 68.10 (0.65) 64.03 (1.06) 83.30 (1.22)
HAF [18] 32.97 (0.57) 68.15 (0.56) 64.26 (0.41) 83.35 (0.68)
HRN [20] 30.48 (0.86) 67.72 (1.98) 63.80 (0.38) 83.05 (0.49)
SAHC (proposed) 34.57 (0.12) 71.25 (0.25) 66.49 (0.34) 84.75 (0.47)
Transformer [36] Baseline (no hierarchy) 18.77 (0.29) 54.79 (0.87) 61.14 (0.62) 81.74 (0.72)
HAF 19.04 (0.28) 55.69 (0.88) 61.12 (0.81) 81.37 (1.03)
HRN 20.97 (0.58) 57.48 (2.31) 60.46 (1.05) 81.80 (0.60)
SAHC (proposed) 25.63 (0.41) 65.94 (0.74) 64.07 (0.45) 83.30 (0.44)
ConvLSTM [38] Baseline (no hierarchy) 25.82 (0.95) 62.59 (0.70) 65.19 (0.77) 84.32 (0.33)
HAF 26.11 (0.92) 62.40 (0.57) 65.64 (0.84) 84.42 (0.68)
HRN 25.53 (1.03) 62.36 (1.47) 65.18 (0.31) 84.19 (0.58)
SAHC (proposed) 26.48 (0.35) 66.74 (0.42) 65.38 (0.40) 85.08 (0.28)
The second study case analyzes the semantic segmentation task, and is characterized by two different datasets with varying hierarchical complexity and differing label quality. The ELU dataset employs weak labels due to its focus on land use rather than land cover. Conversely, the HRLC-CCI dataset is designed for pixel-based LC classification, providing more precise and location-specific labels. Table II shows the performance of the three considered backbones on both the ELU and the HRLC-CCI dataset in terms of mF1% and OA%. The proposed approach outperforms the other methods on the ELU dataset, which exhibits stronger hierarchy and weaker labels than the HRLC-CCI dataset. Notably, the pixel-based transformer without contextual information shows an improvement of about $7\%$ in terms of mF1% compared to the non-hierarchical baseline, while it also outperforms the other hierarchical approaches. In the case of the hybrid Swin Transformer and the ConvLSTM, the performance gap tends to decrease, while the adoption of the hierarchy and the consensus approach still shows an improvement. This result underlines the robustness of the proposed method to noisy labels, as the trained hierarchical matrices can learn soft hierarchical relationships directly from the data. In contrast, the compared methodologies rely on a hard hierarchy definitions, which might lead to suboptimal results in the case of labels that do not accurately represent the observed samples. The same behavior can be observed when analyzing the results on the HRLC-CCI dataset, where the hierarchy becomes weaker, but the real targets are better represented by the labels at the pixel level. In this case, the proposed approach performs similarly to the other methodologies, as the hierarchy is less noisy and the mapping more robust. Nonetheless, SAHC still outperforms the other methodologies, showing good generalization across the different backbones and scenarios. In our experiments, SAHC performed worse than HAF in terms of mF1% score only in the case of the ConvLSTM backbone, although the performance remains comparable. However, the proposed approach still shows a higher OA%, with reduced scores variability across different seeds. Lastly, Fig. 4 shows a visual comparison of the different hierarchical methodologies considering the HRLC-CCI dataset and the Swin Transformer backbone. From the qualitative results, one can see that the SAHC approach better represents the “Tree” classes, also showing a lower pixel-level noise.
Figure 4: Example of qualitative results at the fine-grained level on the HRLC-CCI dataset obtained by the considered methods using the Swin Transformer backbone.
V-C Ablation Study
TABLE III: Ablation study on the Swin backbone with the ELU dataset. The classification metrics are reported with the related standard deviation in brackets. The best results are reported in bold, second best are underlined.
Multi-level
Classifiers Hierarchical
Consensus
(Training) Hierarchical
Consensus
(Inference) mF1 (%)
OA (%)
33.42 (0.44) 68.10 (0.65)
33.48 (0.59) 68.10 (1.06)
34.10 (0.41) 67.83 (1.00)
33.48 (0.37) 68.39 (0.33)
34.57 (0.12) 71.25 (0.25)
The performance of each component in the proposed hierarchical approach is further analyzed in the ablation study presented in Table V-C. In this study, we evaluate the Swin backbone across five distinct configurations on the ELU dataset:
  1. 1.
    Baseline: the baseline approach with no multi-level classifiers;
  2. 2.
    Multi-level: the baseline approach augmented with additional classification heads for the coarse-grained and intermediate classes;
  3. 3.
    Multi-level with HC inference: the multi-level approach with consensus applied solely during inference (considering user-defined hierarchies);
  4. 4.
    Multi-level with $\mathcal{L}_{HC}$ training: the multi-level approach trained with the self-supervised consensus loss, excluding consensus-based classification during inference;
  5. 5.
    SAHC: the proposed approach, which incorporates both the self-supervised consensus loss and consensus inference at multiple levels.
Figure 5: Heatmap of the estimated hierarchy log-joint matrices considering the Swin Transformer backbone on the ELU dataset from fine-grained classes to intermediate classes. The blue and green rectangles identify the expected hierarchical aggregation of the fine-grained classes at the intermediate and coarse levels, respectively.
The table results indicate that incorporating multi-level classifiers into the baseline does not improve the overall performance. In fact, it leads to increased variability in the results, showing a higher variance and making the outcomes more unstable. However, the introduction of the HC inference improves the representation of fine-grained classes in terms of mF1%, also when considering the user-defined hierarchy in the multi-level approach with HC inference. On the contrary, the multi-level approach with $\mathcal{L}_{HC}$ training does not improve the performance, as the JSD acts on the consensus across the layers and not on the fine-grained classification itself. This is instead addressed by the proposed SAHC, which combines the self-supervised consensus loss with consensus inference at multiple hierarchy levels. By allowing the trained matrices to learn the relationships from the data, the hierarchical consensus exhibits a sharp improvement in fine-grained classification, outperforming all other configurations. It is also important to note that the introduction of the JSD and the hierarchical consensus in inference leads to lower variance values, showing a more consistent fine-grained classification across different weights initializations compared to the other approaches.
V-D Hierarchy Projection Matrices Analysis
Figure 6: Heatmap of the estimated hierarchy log-joint matrices considering the Swin Transformer backbone on the HRLC-CCI dataset from fine-grained classes to intermediate classes. The blue and green rectangles identify the expected hierarchical aggregation of the fine-grained classes at the intermediate and coarse levels, respectively.
To better visualize the impact of the hierarchy projection matrices on classification results, we analyze the estimated hierarchy projection matrices of the Swin Transformer on the ELU and HRLC-CCI datasets. The visualization on the NWPU dataset is not reported as the hierarchical mappings show a trivial result, with perfect mapping of the defined hierarchical levels (mostly due to the simpler classification task).
Figure 5 presents a heatmap of the log-joint matrix between intermediate and fine-grained levels, highlighting the hierarchical structure captured by the proposed methodology on the ELU dataset. Note that the hierarchy projection matrix was uniformly initialized before training, with no prior knowledge of the user-defined hierarchical structure of the problem. The results show that the proposed approach effectively learns the hierarchy in the dataset. Specifically, the blue rectangles in the heatmap correspond to the user-defined expected mappings, from fine-grained classes to intermediate classes. Here, the weights associated to these mappings are highlighted, confirming that the data-driven estimated mappings are correctly following the user-defined hierarchies. Then, green rectangles indicate distant mappings to coarse-grained classes, revealing a similar pattern even though the matrix focuses on mappings between fine-grained and intermediate levels. It is important to note that separate hierarchy joint matrices are trained for the different hierarchical mappings in the proposed approach.
Similarly, Figure 6 depicts the heatmap associated with the HRLC-CCI dataset. As observed in the ELU case, the proposed methodology correctly extracts the hierarchy within the problem analyzed, emphasized by the blue rectangle showing the intermediate aggregation. However, the mappings appear more dispersed compared to the more complex ELU scenario, mostly due to the simpler hierarchical constraint leading to a smaller impact. Nonetheless, the trainable log-joint matrices allow soft-hierarchical mappings that were not considered in the initial user-defined hierarchy. One example can be observed in the weights associated to “Grassland vegetation aquatic or regularly flooded areas”, which shows an active semantic relationship with the “Open water” class, which is not mapped by the imposed user-defined hierarchy. This mapping highlights the importance of a data-driven approach, in contrast with the superimposition of a rigid hierarchical structures.
VI Conclusion
In scene classification and semantic segmentation, the hierarchical relationship among the target classes is often under-exploited, resulting in the exclusion of valuable semantic information from the decision-making process. To address this limitation, this paper proposes the SAHC approach that leverages the hierarchical structure inherent in class types to improve the robustness and the accuracy of the classification. We introduced hierarchical classification branches, each designed to handle a specific hierarchical classification task, ranging from coarse-grained to fine-grained categories. Trainable hierarchy log-joint matrices are then defined to project the predictions across different hierarchical levels, following a self-supervised data-driven approach which may not align with the rigid user-defined mappings between different hierarchy levels. This allows the implementation of a consensus mechanism, where predictions from multiple hierarchical levels are fused to perform the final decision. By progressively guiding the classification process from broader to specific categories, we can optimize the consensus, ensuring alignment among predictions at different hierarchical levels in a self-supervised fashion. The results show that data-driven hierarchies improve the performance at the finest level compared to directly using the user-defined hierarchies, contrarily to the performance drop usually observed in the literature when adding additional classification layers for different hierarchical levels after the backbone. The proposed methodology is both lightweight and scalable, making it easily adaptable to different architectures and datasets by simply reorganizing their structure in a hierarchical manner.
We evaluated the performance of the proposed approach on three different RS datasets. Indeed, the proposed approach is particularly relevant in VHR image and RS MS SITS classification, where the semantic structure of the data and available datasets support an explicit hierarchical representation of the task. The experimental results reveal that the proposed methodology is able to exploit both complex and simple hierarchies, demonstrating the best performances in terms of OA% and mF1% scores across experiments at the fine-grained hierarchical level. As expected, the improvements are particularly significant when complex hierarchies can be exploited within the RS data. The proposed method is general, showing top performance scores across different tasks and architectures.
A limitation of the proposed method is the reliance on a predefined hierarchy within the labels to train the self-supervised hierarchical consensus. While the consensus itself operates in a self-supervised manner, its training remains implicitly influenced by the supervised CCE losses at the coarser levels of the hierarchy. This predefined structure may not accurately represent the true underlying relationships within the data, potentially impairing model performance, or constraining the hierarchy log-joint matrices to learn a suboptimal representation of the hierarchical structure. To address this limitation, future research will explore methods for learning the hierarchical structure directly from the data. This approach would eliminate the need for manual definition of hierarchies, thereby improving the adaptability and accuracy of the model.

### v2

@@
-A Semantics-Aware Hierarchical Self-Supervised Approach to Classification of Remote Sensing Images
+A Hierarchical Self-Consistent Regularization Approach to Satellite Image Time Series Classification
 Abstract
@@
 However, these hierarchies are frequently overlooked, and most approaches focus only on fine-grained classification schemes.
-In this paper, we present a novel Semantics-Aware Hierarchical Consensus (SAHC) method for learning hierarchical features and relationships by integrating hierarchy-specific classification heads within a deep network architecture, each specialized in different degrees of class granularity.
-The proposed approach employs trainable hierarchy matrices, which guide the network through the learning of the hierarchical structure in a self-supervised manner.
-Furthermore, we introduce a hierarchical consensus mechanism to ensure consistent probability distributions across different hierarchical levels.
+In this paper, we present a novel Semantics-Aware Hierarchical Consensus (SAHC) approach to learn hierarchical features and relationships by integrating hierarchy-specific classification heads within a deep network architecture, each specialized in different degrees of class granularity.
+The proposed approach employs trainable hierarchy matrices, which guide the network through the learning of the hierarchical structure in a self-consistent manner.
+Furthermore, we introduce a hierarchical consensus mechanism to ensure aligned probability distributions across different hierarchical levels.
 This mechanism acts as a weighted ensemble being able to effectively leverage the inherent structure of the hierarchical classification task.
-The proposed SAHC method is evaluated on three benchmark datasets with different degrees of hierarchical complexity on different tasks, using distinct backbone architectures to effectively emphasize its adaptability.
+The proposed SAHC method is evaluated on two benchmark datasets with different degrees of hierarchical complexity on different tasks, considering varying spectral and spatial resolutions.
 Experimental results show both the effectiveness of the proposed approach in guiding network learning and the robustness of the hierarchical consensus for remote sensing image classification tasks.
+The codes will be released at https://github.com/rslab-unitrento/sahc.
 I Introduction
-Image classification and semantic segmentation are fundamental tasks in image processing.
-While Deep Learning (DL) methods have significantly advanced the field by learning powerful feature representations, standard approaches often treat classification as a flat, single-level task.
-However, many applications involve target categories having a natural hierarchical structure, exhibiting multiple levels of semantic granularity (e.g., object taxonomies, language, land-cover systems) [1, 2, 3].
-Neglecting this inherent structure discards valuable semantic information and can lead to suboptimal or inconsistent predictions [4].
-This is particularly relevant when dealing with label ambiguity or variations in image quality (such as resolution or artifacts) that may only permit reliable classification at coarser semantic levels.
-A key challenge is integrating hierarchical information effectively into the learning process.
-Traditional methods frequently rely on post-processing steps or fixed rule-based systems to enforce hierarchical consistency, which limit their ability to learn complex, data-driven relationships between levels [5, 6, 7, 8].
-Furthermore, constructing large-scale datasets with consistent multi-level annotations is often difficult [9, 10].
-This requires methods capable of handling data labelled at various levels within the hierarchy.
-Although standardized hierarchical classification schemes exist in several domains, e.g., COoRdination of INformation on the Environment (CORINE) and Land Cover Classification System (LCCS) in Remote Sensing (RS), most current DL models do not explicitly leverage this structure during training.
-Most hierarchical approaches focus on extracting features at multiple levels of granularity without explicitly modeling the hierarchical relationships intrinsic to the task [11, 12].
-This poses a limitation, as the hierarchy arises from the problem domain itself rather than from specific architectural design choices.
-Accordingly, an effective hierarchical learning principle should be generalizable and scalable across different DL architectures.
-However, existing methods often produce non-modular architectures with rigid structures, thereby limiting the flexibility in the DL backbone selection.
-In this paper, we propose a novel image processing methodology to modularly improve scene classification and semantic segmentation through semantic relationships across categories.
-Our approach leverages the hierarchical structure of standardized classification schemes by introducing a classification head for each hierarchical level.
-We further introduce trainable hierarchy matrices to model the relationships between these levels.
-To enforce consistency across the hierarchy, we define a multi-level consensus prediction that promotes agreement among the predictions at different levels.
-Additionally, we introduce a self-supervised consensus loss that implicitly learns the hierarchical relationships within the data.
-This approach allows the discovery of intricate sematic relationships that can deviate from the simple user-defined aggregation rules, leading to improved classification accuracy and semantic consistency.
-We demonstrate the effectiveness of our approach on different RS image classification tasks, including semantic segmentation of Satellite Image Time Series (SITS) of multispectral acquisitions and Very-High Resolution (VHR) RS scene classification tasks.
+Image classification and semantic segmentation are central tasks in Remote Sensing (RS) image processing.
+In recent years, the proliferation of RS data and open distribution policies have significantly increased the availability of Land-Use and Land-Cover (LULC) products, along with their corresponding classification schemes [27].
+These products play a critical role in various applications and are widely used by government agencies, municipalities, and ministries.
+However, global and regional LULC maps often differ in terms of resolution and are affected by semantic inconsistencies in class definitions and production algorithms [28].
+While Deep Learning (DL) methods have revolutionized LULC mapping by capturing complex spatial and spectral dependencies, standard approaches typically treat classification as a flat, single-level task.
+In practice, Land-Cover (LC) classes are naturally organized into hierarchical label structures that reflect intrinsic levels of semantic granularity.
+This multi-level hierarchy is a property of the physical landscape.
+For instance, within a single dataset, a specific target might be unambiguously identified at a coarse-grained level (e.g., forest), while its finer-grained classification (e.g., deciduous versus coniferous) remains uncertain due to inherent label ambiguity, complex scene compositions, or the physical limits of the sensor.
+Neglecting this inherent hierarchical structure discards valuable semantic information and can lead to suboptimal or inconsistent predictions, particularly when dealing with label ambiguity or complex RS scenes.
+An example of this hierarchical problem definition can be seen in Fig.
+1.
+Figure 1: Hierarchical classification task example. Level 4 represents the coarsest class definition, where each class is progressively broken down into its fine-grained representations.
+In the literature, most of the works related to RS image classification focus on single-level approaches, discarding the hierarchical information contained in the data.
+Hierarchical classification schemes, such as the CORINE Land Cover (CLC) maps, have mostly been adopted in object-based image classification methods based on Very-High Resolution (VHR) satellite images [3], often considering a limited number of classes to simplify the classification task.
+To validate the hierarchical structure, few works analyzed the inter-level relationships, considering decision trees in an object-based classification scenarios [25], or validation approaches based on intercomparison of different maps [23].
+Instead, the wide majority of works in the RS domain focus on single-level approaches, discarding the hierarchical information contained in the data.
+While these methods show promising results at individual hierarchical levels, they fail to incorporate the hierarchical structure of the data into the decision-making process, overlooking important information included in the source data.
+Only recently, hierarchical and multi-granularity classification methods have emerged to address these challenges in RS.
+These methods leverage the hierarchical structure to encode semantic relationships among labels, enabling knowledge transfer across different levels of granularity.
+Recent studies have demonstrated the benefits of adaptive label granularity selection to balance classification accuracy and specificity [7], as well as the integration of parent-child relationships and sibling mutual exclusivity to improve hierarchical representation learning [9].
+Despite these advancements, integrating hierarchical information effectively into the learning process remains an open challenge.
+Many existing methods rely on complex, task-specific architectures or rigid rule-based post-processing steps, which limit their flexibility across different DL backbones and fail to explicitly enforce probability consistency across different hierarchical levels in a generalized manner.
+To address these limitations, we propose a novel and modular Semantics-Aware Hierarchical Consensus method for learning hierarchical features and relationships.
+Our approach integrates hierarchy-specific classification heads within standard DL architectures, each specialized in different degrees of class granularity.
+By introducing trainable projection matrices and explicitly modeling the semantic relationships between these levels, the proposed method ensures consistent predictions across the hierarchy without relying on rigid aggregation rules.
 The key contributions of this work can be summarized as follows:
   1. 1.
-    We define a hierarchical image processing approach to semantic segmentation and scene classification that integrates multi-level classification modules with DL backbones, along with trainable projection matrices to model hierarchical semantic relationships.
+    We define a hierarchical approach to semantic segmentation and scene classification that integrates multi-level classification modules with DL backbones, along with trainable projection matrices to model hierarchical semantic relationships.
   2. 2.
-    We propose a novel self-supervised consensus loss function that enforces consistency across different hierarchical levels, enhancing the semantic coherence of the classification results.
+    We propose a novel self-consistent consensus loss function that enforces consistency across different hierarchical levels, improving the semantic coherence of the classification results.
   3. 3.
-    The proposed method is general and can be applied to any DL architecture and hierarchical classification scheme, offering flexibility in defining architecture details.
+    The proposed method is general and can be applied to any DL backbone and hierarchical classification scheme, offering flexibility in defining architecture details.
   4. 4.
-    We demonstrate the effectiveness of the proposed approach on three RS datasets that include images with varying resolutions and different complexities in their hierarchical relationships.
+    We demonstrate the effectiveness of the proposed approach on two RS datasets that include images with varying resolutions and different complexities in their hierarchical relationships.
 The remainder of this paper is organized as follows.
@@
 II Related Works
-Image classification is a fundamental task in computer vision and image processing, with DL models achieving state-of-the-art performance across various domains.
-Standard approaches typically treat classification as a “flat” problem in a traditional single-granularity classification, predicting probabilities over a set of mutually exclusive classes without explicitly considering potential semantic relationships or hierarchical structures among them.
-However, many real-world datasets possess inherent taxonomies where classes can be organized into hierarchies (e.g., species classification, object recognition with super-/sub-categories).
-Ignoring this structure can lead to suboptimal performance, as the model is not able to leverage prior knowledge about class relationships.
-Consequently, standard loss functions inherently treat all misclassifications equally.
-Without an explicit hierarchical representation, they fail to differentiate between errors involving closely related classes and those involving semantically distant ones.
+Image classification and semantic segmentation are widely studied tasks in RS image processing, with DL models achieving state-of-the-art performance across various domains.
+Standard approaches typically treat classification as a “flat”, single-granularity problem, predicting probabilities over a set of mutually exclusive classes without explicitly considering potential semantic relationships or hierarchical structures among them.
+However, many real-world RS datasets possess inherent taxonomies where classes can be naturally organized into hierarchies (e.g., standardized land-cover systems, target recognition taxonomies with super- and sub-categories).
+Ignoring this structure can lead to suboptimal performance, as standard loss functions treat all misclassifications equally.
+Without an explicit hierarchical representation, model fail to differentiate between errors involving closely related classes and those involving semantically distant ones.
 Leveraging hierarchical label structures during model training and inference has emerged as a promising direction to improve accuracy and semantic understanding.
@@
 Building on this, researchers have explored explicit methods to incorporate hierarchy.
-Typically, this is achieved by mapping the label hierarchy to network architectures or by introducing loss functions that handle the hierarchical correlation between different granularity tasks [14, 15, 16].
+Typically, this is achieved by mapping the label hierarchy to specific network architectures or by introducing loss functions that handle the correlation between different granularity tasks [5, 14, 17].
+For instance, several methods propose level-wise hierarchical cross-entropy losses based on conditional probabilities, or introduce custom features to minimize the divergence between aggregated predictions at different hierarchical levels, aiming to reduce the severity of misclassification.
 Bertinetto et al.
@@
 They considered exclusions, overlaps, and subsumptions, and provided a detailed theoretical analysis of the structure of the hierarchical problem.
-Building on this work, a combinatorial loss is proposed in [20], which addresses the scenario where samples might be labelled at different levels of granularity.
+Building on this work, a combinatorial loss is proposed by Chen et al.
+[8], which addresses the scenario where samples might be labelled at different levels of granularity.
 The loss is designed to maximize the marginal probability of the observed ground truth label by aggregating information from related labels in the hierarchy.
 The authors introduced a Hierarchical Residual Network (HRN) architecture, where features from parent levels act as residual connections to inform child-level features, explicitly facilitating hierarchical feature interaction.
-While these methods demonstrate the benefits of hierarchy-aware learning, they often rely on the assumption that a predefined and accurate hierarchy is available, where no hierarchy violation is possible.
-This assumption may not hold in all scenarios, as real-world taxonomies can be noisy, incomplete, or subjective.
+Following this work, Zhou et al.
+[31] proposed a multi-granularity archaeological dating method that depends on an explicit knowledge-guided relation graph, employing a focal-type probability classification loss to mine the relationship between specific object attributes and their broader eras.
+Their network leverages a bidirectional multi-granularity module with gradient truncated addition to explicitly fuse coarse- and fine-grained features.
+Similarly, Cai et al.
+[4] introduced HierSwin, a hierarchical vision transformer, which explicitly models the semantic correlations embedded within a predefined three-level hierarchical tree to gradually improve accuracy from fine- to coarse-grained classifications.
+More recently, some works addressed the Multigranularity Classification (MGC) in RS to handle targets that are discernible only at specific variable semantic levels due to differences in spatial resolution, sensor modalities, and data quality.
+Chen et al.
+[7] introduced an accuracy-specificity curve to adaptively select task-dependent prediction thresholds.
+Their approach also employs a penalty term based on the global semantic distance across the label tree to adjust class logits.
+Addressing a similar challenge in ship classification, Chen et al.
+[9] proposed a hierarchical contrastive learning framework.
+By explicitly incorporating parent-child dependencies and mutual exclusivity among sibling categories, their method aligns class-specific representations with semantic relationships and enforces probability coherence across levels through a consistency loss.
+Despite these significant advancements, many existing multigranularity and hierarchical learning methods still rely on the assumption that a predefined and accurate hierarchy is available, where no hierarchy violation is possible.
+This requires rigid post-processing rules, fixed decision thresholds, or highly specific non-modular architectures.
 Furthermore, many existing techniques require explicit guidance or architectural modifications tailored to a specific fixed hierarchy [18], potentially limiting their flexibility and ability to adapt if the underlying hierarchical understanding evolves or is uncertain.
-For example, some approaches decompose the problem into sequential decisions or binary classifications guided by the known structure [22, 23, 21], rather than allowing the model to learn or refine hierarchical relationships implicitly from the data.
+For example, some approaches decompose the problem into sequential decisions or binary classifications guided by the known structure [11, 29], rather than allowing the model to learn or refine hierarchical relationships implicitly from the data.
 The challenge remains to develop methods that can effectively exploit hierarchical semantics, even when the hierarchy is imperfect, labels are provided at varying granularity levels, or needs to be partially inferred in a robust and data-driven manner.
@@
 This allows the model to leverage potential hierarchical structures effectively while mitigating the impact of potential inaccuracies or inconsistencies in the predefined taxonomy.
-Specifically, our method aims to learn and exploit hierarchical relationships in a more implicit, self-supervised fashion within a hierarchy-aware architecture, reducing the reliance on rigid, externally defined structure.
+Specifically, our method aims to learn and exploit hierarchical relationships in a more implicit, self-consistent fashion within a hierarchy-aware architecture, reducing the reliance on rigid, externally defined structure.
 III-A Hierarchical Problem Formulation
@@
 Its associated labels are $y^{h}\in\Omega^{h}$, where $\Omega^{h}=\{\omega^{h}_{1},\dots,\omega^{h}_{|\Omega^{h}|}\}$ is the set of labels at hierarchical level $h$, with $|\Omega^{h}|$ number of labels.
-For every $h_{1}<h_{2}$, we have that each fine-grained class $\omega_{j}^{h_{2}}$ is mapped to a coarse-grained class $\omega_{i}^{h_{1}}$, and $|\Omega^{h_{1}}|<|\Omega^{h_{2}}|$.
+For every $h_{1}<h_{2}$, we have that each fine-grained class $\omega_{j}^{h_{2}}$ is mapped to a coarse-grained class $\omega_{i}^{h_{1}}$, and $|\Omega^{h_{1}}|\leq|\Omega^{h_{2}}|$.
 We define this mapping relationship by $\omega_{j}^{h_{2}}\implies\omega_{i}^{h_{1}}$.
@@
 2 provides a high-level overview of the considered multi-level classification approach.
-Figure 1: Overview of the proposed hierarchical structure within the backbone (left) and associated hierarchical label tree (right).
-Different classification modules are trained to classify different hierarchical levels from the features extracted by the backbone.
+Figure 2: Illustration of the proposed hierarchical architecture with the hierarchical consensus calculated on $h=\tilde{h}$.
+Heads tailored to different granularity levels are reprojected on the target level and averaged to extract the consensus probability $\hat{p}^{\tilde{h}}_{HC}$.
+Other than for prediction, this consensus probability is used to calculate the hierarchical self-consistency loss $\mathcal{L}^{h}_{HC}$.
 Each classifier $\mathbf{g}^{h}$ estimates class-posterior logits.
@@
 The CCE loss functions at each hierarchical level lead to independent classifications, which are not constrained to follow the hierarchy definition formulated in the problem statement.
-To impose the hierarchical constraints, additional consensus predictions and self-supervised penalty terms are introduced.
+To impose the hierarchical constraints, additional consensus predictions and self-consistent penalty terms are introduced.
 These penalty terms are defined through hierarchy projection matrices designed to map the correct relationships between the different layers.
@@
 We refer to $\Omega^{h_{1}}$ as the set of the macroclasses (i.e., the coarse-grained classes) and to $\Omega^{h_{2}}$ as the set of the microclasses (i.e., the fine-grained classes).
-We can define the hierarchy indicator matrix $\mathbf{I}^{h_{2},h_{1}}\in\{0,1\}^{|\Omega^{h_{2}}|\times|\Omega^{h_{1}}|}$, which represents the hierarchical relationship between level $h_{2}$ and $h_{1}$:
+We can define the hierarchy indicator matrix $\mathbf{I}^{h_{2},h_{1}}\in\{0,1\}^{|\Omega^{h_{2}}|\times|\Omega^{h_{1}}|}$, which represents the user-defined hierarchical relationship between level $h_{2}$ and $h_{1}$:
 $\mathbf{I}^{h_{2},h_{1}}=\begin{bmatrix}I_{1,1}&I_{1,2}&\dots&I_{1,|\Omega^{1}|}\\ I_{2,1}&I_{2,2}&\dots&I_{2,|\Omega^{1}|}\\ \vdots&\vdots&\ddots&\vdots\\ I_{|\Omega^{2}|,1}&I_{|\Omega^{2}|,2}&\dots&I_{|\Omega^{2}|,|\Omega^{1}|}\end{bmatrix}$.
@@
 In this case, the total number of hierarchy matrices that can be determined is $((H-1)\cdot H)/2$, as also mappings of distant relationship levels can be defined (e.g., from the root of the label tree to the leaves, skipping intermediate levels) if needed.
-An example is illustrated in Fig.
-2.
+A generalized example is illustrated in Fig.
+3.
 Figure 3: Illustration of both direct and distant hierarchical mappings, from fine-grained classes to coarse-grained classes.
-The hierarchy indicator matrices are initialized with ones and zeros representing the user-defined mapping.
-Consequently, $I^{h_{2},h_{1}}_{i,j}$ assumes a value of one when the microclass $\omega_{i}^{h_{2}}$ is contained in the macroclass $\omega_{j}^{h_{1}}$ through the relationship $\omega_{i}^{h_{2}}\implies\omega_{j}^{h_{1}}$}, thus defining the matrix elements as follows:
-$I^{h_{2},h_{1}}_{i,j}=\begin{cases}1&\text{if }{\omega_{i}^{h_{2}}\implies\omega_{j}^{h_{1}}}\\ 0&\text{otherwise}\end{cases}.$ (6)
-It is important to note that, in the case of hard-hierarchies, each microclass corresponds to exactly one macroclass.
-However, this may not hold true as there might exist semantic overlaps in the hierarchy structure.
-Additionally, a distinction between user-defined aggregations and data-driven aggregations has to be made.
-While the first is defined a-priori for the problem at hand by the user, the latter is not known and must be inferred directly from the data, reflecting the true underlying structure of the problem considered.
-To take this into account, we can soften the hierarchical relationships by allowing the matrices to take any value between 0 and $1$.
-In this way, one microclass may be associated with multiple macroclasses with different confidences reflecting the semantic overlaps.
-To this purpose, we can define hierarchy joint distribution matrices $\mathbf{J}^{h_{2},h_{1}}\in{[0,1]}^{|\Omega^{h_{2}}|\times|\Omega^{h_{1}}|}$.
+In a strict hard-hierarchy, $I^{h_{2},h_{1}}_{i,j}$ assumes a value of one when the microclass $\omega_{i}^{h_{2}}$ is contained in the macroclass $\omega_{j}^{h_{1}}$ through the relationship $\omega_{i}^{h_{2}}\implies\omega_{j}^{h_{1}}$}, and zero otherwise.
+However, this rigid formulation often fails to account for semantic overlaps in real-world hierarchies.
+Therefore, we distinguish between user-defined aggregations, established a priori for a given task, and data-driven aggregations, which are inferred directly to capture the data’s true underlying organization.
+To combine the user-defined prior with a data-driven optimization, we parameterize the hierarchical relationships in the unconstrained logit space.
+Instead of stricly fixing the matrices to binary values or directly bounding them in $[0,1]$, we define learnable parameter matrices $\mathbf{W}^{h_{2},h_{1}}$.
+These matrices are initialized based on the user-defined indicator matrix $\mathbf{I}^{h_{2},h_{1}}$ to provide a stronger prior:
+$W^{h_{2},h_{1}}_{i,j}=\begin{cases}0.5+\epsilon&\text{if }{I^{h{2},h_{1}}_{i,j}=1}\\ -5.0+\epsilon&\text{if }{I^{h{2},h_{1}}_{i,j}=0}\end{cases},$ (6)
+where $\epsilon\sim\mathcal{N}(0,0.01^{2})$ is a small random noise added to break network symmetry during initial phases of optimization.
+The two selected values act as positive and negative biases towards valid connections and unmapped relationships.
+During the forward pass, these unconstrained parameters are dynamically mapped into valid hierarchy log joint distribution matrices $\mathbf{L}^{h_{2},h_{1}}=\log{\mathbf{J}^{h_{2},h_{1}}}\in{\mathbb{R}}^{|\Omega^{h_{2}}|\times|\Omega^{h_{1}}|}$ via a log-softmax operation on the flattened parameter matrices $\mathbf{W}^{h_{2},h_{1}}$.
 Each term of the matrix $J^{h_{2},h_{1}}_{i,j}=P(\omega^{h_{1}}_{j},\omega^{h_{2}}_{i})$ models the probability of observing a relationship between microclass $\omega^{h_{2}}$ and macroclass $\omega^{h_{1}}$.
-This constraint relaxation allows us to better combine the user-defined aggregation problem with a data-driven centered architecture, by learning these matrices directly from the data.
-III-D Self-Supervised Hierarchical Consensus Learning
+This constraint relaxation allows us to better combine the user-defined aggregation with a data-driven centered architecture, by learning these matrices directly from the data while still reliably considering the initial user-defined mappings.
+III-D Self-Consistent Hierarchical Consensus Learning
 Let us return to the general case of $H$ hierarchical levels, considering projections from source levels $h$ to a target hierarchical level $\tilde{h}$.
@@
 All predictions and their projected versions can be viewed as “soft” votes, which define each committee member’s confidence.
-Then, we introduce a self-supervised loss that penalizes the committee members’ disagreement with the overall committee consensus, thus guiding the learning of data-driven relationships in $\mathbf{J}^{h,\tilde{h}}$ and providing further supervision to the network based on the predictions at the different levels.
+Then, we introduce a self-consistent loss that penalizes the committee members’ disagreement with the overall committee consensus, thus guiding the learning of data-driven relationships in $\mathbf{J}^{h,\tilde{h}}$ and providing further supervision to the network based on the predictions at the different levels.
 III-D1 Hierarchy-based Projected Predictions
@@
 $\log{\left(P^{h\rightarrow\tilde{h}}\left(\omega_{j}^{\tilde{h}}|\mathbf{x}\right)\right)}=\\ =\text{LSE}\left(\mathbf{g}^{h}+\log{\left(\mathbf{\hat{M}}^{h,\tilde{h}}_{\cdot,j}\right)}\right)-\text{LSE}\left(\mathbf{g}^{h}\right).$ (12)
-To further simplify, hierarchy projection matrices can be defined directly in the log domain, i.e., we learn log joint probabilities $\mathbf{L}^{h,\tilde{h}}=\log{\hat{\mathbf{J}}^{h,\tilde{h}}}\in\mathbb{R}^{|\Omega^{h}|\times|\Omega^{\tilde{h}}|}$.
+Hierarchy projection matrices can be defined directly in the log domain, i.e., we learn log joint probabilities $\mathbf{L}^{h,\tilde{h}}=\log{\hat{\mathbf{J}}^{h,\tilde{h}}}\in\mathbb{R}^{|\Omega^{h}|\times|\Omega^{\tilde{h}}|}$.
 Then, we compute $\log{\hat{\mathbf{M}}^{h,\tilde{h}}}$ as follows:
@@
 where $\text{LSE}{(\mathbf{g}^{\tilde{h}}_{HC})}=\log{Z^{\tilde{h}}}$.
-While not directly supervised, HC can serve as an in-place replacement for the original classification modules during inference, providing a fully semantics-aware prediction in the downstream task.
-III-D3 Self-Supervised Consensus Loss
+This HC represents the consensus found by the voting committee and can serve as an in-place replacement for the original classification modules during inference, providing a fully semantics-aware prediction in the downstream task.
+III-D3 Self-Consistent Consensus Loss
 After calculating the consensus, both the prediction at the target level and the projected predictions are compared with the HC, defining an additional loss based on the divergence from the HC.
-To this purpose, we employ the JSD, which is defined as follows:
+To this purpose, we emply a JSD loss rather than a standard Kullback–Leibler divergence (KLD).
+While the KLD is widely used to measure the difference between two probability distributions, its asymmetry makes it better suited for scenarios where one distribution is a fixed “ground truth” or prior.
+In this case, all hierarchical levels provide dynamically trained “soft” votes, with no reference distribution.
+The symmetry in the JSD ensures balanced regularization between the projected predictions and the consensus, also providing stabler gradients due to its bounded nature.
+In the proposed framework, the JSD loss is computed as follows:
 $\text{JSD}^{h,\tilde{h}}\left(\hat{\mathbf{p}}^{\tilde{h}}_{HC}||\hat{\mathbf{p}}^{h\rightarrow\tilde{h}}\right)=\\ =\frac{1}{2}\left(\text{KLD}\left(\hat{\mathbf{p}}^{\tilde{h}}_{HC}||\mathbf{q}\right)+\text{KLD}\left(\hat{\mathbf{p}}^{h\rightarrow\tilde{h}}||\mathbf{q}\right)\right),$ (16)
@@
 […]
+Finally, we also apply full supervision on the consensus predictions at the different levels considering a standard CCE loss:
+$\mathcal{L}_{CCE}^{HC}=\sum_{h=1}^{H}-\frac{1}{N}\sum_{n=1}^{N}\sum_{i=1}^{|\Omega^{h}|}\mathbbm{1}\left[y^{h}_{n}=\omega^{h}_{i}\right]\log\hat{P}^{h}_{HC}\left(\omega^{h}_{i}|\mathbf{x}_{n}\right).$ (18)
 III-E Overall Loss Function
 Lastly, the losses are averaged and summed to $\mathcal{L}_{CCE}$ (Eq. 4):
-$\mathcal{L}=\mathcal{L}_{CCE}+\lambda_{HC}\mathcal{L}_{HC},$ (18)
+$\mathcal{L}=\mathcal{L}_{CCE}+\mathcal{L}_{CCE}^{HC}+\lambda_{HC}\mathcal{L}_{HC},$ (19)
 where $\lambda_{HC}$ is a metaparameters tuning the contribution of the HC loss.
-It is important to note that $\mathcal{L}_{HC}$ does not require supervision, as it is simply a loss representing the divergence from the HC.
-Though the HC prediction is used within the penalty loss during the training phase, we consider the HC prediction at each hierarchical level as the final predictions for inference.
-This allows the architecture to leverage the semantic relationships learnt during the self-supervised training phase and to perform hierarchically consistent predictions.
+It is important to note that $\mathcal{L}_{HC}$ does not require direct supervision, as it is simply a loss representing the divergence from the HC.
+The HC prediction is used both within the penalty loss, and the additional supervised CCE loss during the training phase.
+We further consider the HC prediction at each hierarchical level as the final predictions for inference.
+This allows the architecture to leverage the semantic relationships learnt during the self-consistent training phase and to perform hierarchically consistent predictions.
 IV Dataset Description and Experiment Setup
@@
 IV-B ELU Dataset
-Figure 3: A semicircular hierarchical label tree representation of the Emilia LU dataset at the fine-grained level of the hierarchy.
-The leaves corresponding to the fine-grained classes are shown in green, while the nodes for intermediate classes are depicted in blue, and those for coarse-grained classes are shown in red.
+Figure 4: A semicircular hierarchical label tree representation of the ELU dataset at the fine-grained level of the hierarchy.
+The leaves corresponding to the fine-grained classes are shown in orange, while the nodes for intermediate classes are depicted in green, those for coarse-grained classes are shown in blue, and the nodes for coarse in brown.
 The root node is represented in black.
-To further validate the performance of the proposed methodology, we additionally considered two multitemporal multispectral datasets differing in terms of complexity in the hierarchical structure.
-The first dataset [^1] Land Use of Emilia (2020).
-Accessed on May $13^{\text{th}}$ $2025$.
-consists of a publicly available LU map of $2020$ over the Emilia Romagna region, Italy.
-The maps were defined considering the standard CORINE Land Cover (CLC) aggregation at three levels, with the third one further divided, resulting in an additional fourth level with an increased semantic detail.
-For the purpose of this analysis, we evaluated the hierarchical approach on three levels of hierarchy, consisting in the first two levels of the CLC and the additional finest level of detail, here referred to as coarse-grained, intermediate and fine-grained, respectively.
+To further validate the performance of the proposed methodology and demonstrate its generability on different land cover mapping tasks, we additionally considered a multitemporal multispectral dataset differing in terms of complexity in the hierarchical structure.
+This dataset [^1] Land Use of Emilia (2020).
+Accessed on April $9^{\text{th}}$ $2026$.
+consists of a publicly available Land-Use (LU) map of $2020$ over the Emilia Romagna region, Italy.
+The maps were defined considering the standard CLC aggregation at four levels, showcasing a highly detailed semantic dataset.
+For the purpose of this analysis, we evaluated the hierarchical approach on four levels of hierarchy, consisting in the first three levels of the CLC and the additional finest level of detail defined by the LULC, here referred to as macro, coarse-grained, intermediate and fine-grained, respectively.
 The dataset has been built considering orthophotos and infrared images, and shows a minimum mapping unit of $0.16$ hectares, with a minimum of $7\text{m}$ meters for linear structures.
-Although it has a good positional accuracy of $5\text{m}$, the data may exhibit errors.
-Additionally, it represents LU classes, where the same Land-Cover (LC) class could be associated with different semantic classes, such as the fine-grained classes “Commercial airports”, “Sport flight airports and heliports”, and “Military airports”.
-We considered the Emilia Land Use (ELU) dataset corresponding to the Military Grid Reference System (MGRS) Sentinel-2 (S2) tile 32TPQ ($10980\times 10980$ pixels, for a total area of $110\times 110$ $km^{2}$).
-In the analyzed tile, the three hierarchical levels show $5$, $18$, and $80$ classes at the coarse-grained, intermediate, and fine-grained levels, respectively.
+Additionally, it represents LU classes, where the same LC class could be associated with different semantic classes, such as the fine-grained classes “Commercial airports”, “Sport flight airports and heliports”, and “Military airports”.
+We considered the ELU dataset corresponding to the Military Grid Reference System (MGRS) Sentinel-2 (S2) granule 32TPQ ($10980\times 10980$ pixels, for a total area of $110\times 110$ $km^{2}$).
+In the analyzed granule, the four hierarchical levels show $5$, $15$, $40$, and $90$ classes at the macro, coarse-grained, intermediate, and fine-grained levels, respectively.
 Please note that such an elevated number of fine-grained classes is not commonly approached in Earth Observation (EO) classification problems, where the majority of studies focus on the intermediate or coarse-grained classification task.
@@
 A semicircular tree representing the hierarchical structure of the dataset is shown in Fig.
-3.
-IV-C ESA HRLC CCI Dataset
-The second multitemporal multispectral dataset [^2] ESA High-Resolution Land Cover Climate Change Initiative Project CEDA Archive site.
-Accessed on May $13^{\text{th}}$ $2025$.
-has been derived under the High Resolution Land Cover (HRLC)- Climate Change Initiative (CCI) project by European Space Agency (ESA) [30].
-The dataset includes LC classes organized according to the LCCS nomenclature, following the classification at the finest level of details.
-In this case, the hierarchical structure is less rigid compared to the Emilia dataset, as the classes are less detailed and can almost be considered independent.
-In this analysis, we use the HRLC1010 maps, i.e., the $10\text{m}$ resolution maps for the year $2019$.
-The maps have been derived from machine learning approaches considering a joint approach based on multispectral (MS) S2 and Synthetic Aperture Radar (SAR) Sentinel-1 (S1) images [31].
-The HRLC1010 maps have been validated by intercomparison with other available maps and with photointerpreted samples.
-For our analysis, we considered the dataset corresponding to the MGRS S2 tile 22KGV ($10980\times 10980$ pixels, for a total area of $110\times 110$ $km^{2}$) in Brazil, mostly due to the heterogeneous presence of different classes.
-Since the hierarchical structure is simpler compared to the Emilia LU dataset, we defined three hierarchical levels ad-hoc, based on the similarity of the classes at the finest level of detail.
-This led to the definition of $3$, 77, and $12$ classes for the different levels, corresponding to coarse-grained, intermediate, and fine-grained classes, respectively.
-The coarse-grained classes have been divided in “Woody vegetated areas”, “Herbaceous vegetated areas”, and “Non-vegetated areas”.
-The fine-grained classes have been aggregated at the intermediate level considering the seasonality (e.g., “Tree cover evergreen broadleaf” and “Tree cover deciduous broadleaf” are aggregated at intermediate level, while “Tree cover evergreen needleleaf” is still separated and aggregated at the coarse-grained level instead).
-Since the number of classes only shows a slight increment in the different layers, we refer to this hierarchy as a simple hierarchy.
-IV-D S2 Preprocessing and Dataset Preparation
-IV-D1 VHR dataset
+4.
+IV-C S2 Preprocessing and Dataset Preparation
+IV-C1 VHR dataset
 The training, validation, and test splits were selected accordingly to [21], considering a split-ratio of $0.6$, $0.2$, and $0.2$, respectively.
 During training, we applied data augmentation techniques consisting in resize, random crops and random rotations of the images.
-IV-D2 Multitemporal MS datasets
-The input data consists of S2 MS satellite images provided in the Copernicus EO program.
+IV-C2 Multitemporal MS datasets
+The input data consists of S2 multispectral (MS) satellite images provided in the Copernicus EO program.
 The S2 acquisitions considered have been preprocessed at Level-2A (L2A) and re-projected at $10\text{m}$/pixel resolution through a nearest neighbor approach.
 In the experiments, all the bands except the $60\text{m}$ resolution band were used.
-Due to the high-complexity of both classification problems, we collected two time series (TS) of S2 images, fully exploiting the temporal signature and allowing the detection of seasonal targets.
-The S2 TS considered for the Emilia Dataset consists of acquisitions ranging from January $1^{\text{st}}$ $2020$ to December $31^{\text{st}}$ $2020$, while the S2 TS for the HRLC-CCI dataset consists of TS ranging from January $1^{\text{st}}$ $2019$ to December $31^{\text{st}}$ $2019$.
-Both the S2 TSs have been processed to achieve more regular TSs, resulting in $12$ median monthly composites.
-This was achieved by computing the pixel-wise median for each band for each month.
-Before applying the median operator, we masked the samples associated with saturate or defective pixels, clouds, and cloud shadows.
-This approach allowed us to have regular, gap-free data in input to the DL models.
-To extract a balanced training set, we considered an approach based on the heterogeneity of the input training patches.
-In particular, for both datasets, we divided the initial S2 monthly composite images into non-overlapping patches of $9\times 9$ pixels.
-After the initial subdivision, we collected the patches and calculated the mode, extracting the dominant class of each patch analyzed.
-Then, the patches have been randomly subdivided into training, validation, and test sets, by considering a minimum number of $N$ representative patches for each class in the analyzed classification scheme, with $N$ fixed to $1000$ for both the ELU and HRLC-CCI datasets.
-In the case of a minoritarian class had fewer patches than $N$, $50\%$ of its total were assigned to the training set, $25\%$ to the validation set, and $25\%$ to the test set.
-In this way, we collected a small training dataset representative of the entirety of the classification scheme, corresponding to $\sim 1.35\%$ ($1.631.616$ pixels) and to $\sim 0.22\%$ ($268.800$ pixels) of the total number of patches in the tile for the Emilia Dataset and the HRLC-CCI dataset, respectively.
-The validation set size has been determined by considering $25\%$ of the number of samples allocated for each class for the training set, while the remainder of the samples (i.e., the entirety of the patches in the tile not belonging to the training and validation sets) have been used to validate the proposed approach in the test set.
-IV-E Design of Experiments
-Instead of validating the proposed hierarchical approach on a single backbone, we demonstrated its broad applicability by employing diverse backbone architectures representing different paradigms in a robust manner.
-All tests were conducted five times, with different random initializations of the weights of the architecture.
-For reproducibility, predefined seed values were used for the random number generator (RNG).
-The reported results include the values of Overall Accuracy (OA) for both tasks, while mean F1 Score (mF1) was also employed in the SITS MS classification task, along with the standard deviation (calculated as an unbiased estimator).
-For each architecture, we considered the following approaches: (i) the baseline with only the fine-grained classification through standard CCE minimization, (ii) the HAF method [18], (iii) the HRN module [20] and (iv) the proposed Semantics-Aware Hierarchical Consensus (SAHC).
-We removed the Granularity-Specific Block (GSB) block (two parameter-heavy convolutions and fully connected layers) from the HRN approach, maintaining only the hierarchical module.
-This is to ensure a fair comparison of architectures with the same number of parameters and focus solely on the hierarchical representation of the problem.
-The different strategies are evaluated on the VHR scene classification task and the multitemporal MS semantic segmentation tasks.
-For the VHR task, we considered three different backbone architectures focusing on three different paradigms: (i) EfficientformerV2 S2 [33] for transformer-based architectures, (ii) MobileNetV2 [34] for CNN-based architectures, and (iii) LWGANet [35] for hybrid architectures.
-For the MS SITS classification task, we considered three different backbone architectures designed to process temporal information, also including hybrid architectures to further examine the effectiveness of the methodologies in utilizing contextual information: (i) the encoder part of a Transformer [36], (ii) an adaptation to the RS scenario of a Swin Transformer originally intended for super-resolution of videos [37], and (iii) a Convolutional Long Short-Term Memory (ConvLSTM) [38].
-For a detailed description of the layer topology, please refer to the cited papers.
-For each backbone, we replaced the classification head with three hierarchical classification heads, each characterized by a different level of granularity.
+Due to the high-complexity of both classification problems, we collected a time series (TS) of S2 images, fully exploiting the temporal signature and allowing the detection of seasonal targets.
+The S2 TS considered for the Emilia Dataset consists of acquisitions ranging from January $1^{\text{st}}$ $2020$ to December $31^{\text{st}}$ $2020$.
+The S2 TS has been processed to achieve a more regular TS, resulting in $12$ medoid monthly composites [13].
+To extract a balanced training set, we considered an approach based on the heterogeneity of the input training data.
+In particular, we divided the initial S2 monthly composite images into non-overlapping tiles of $1000\times 1000$ pixels ($10km^{2}$).
+After the initial subdivision, we randomly subdivided the tiles into training, validation, and test sets by considering a split-ratio of $0.7$, $0.15$ and $0.15$, respectively.
+We extracted $N=2000$ representative points for each tile and class in the analyzed classification scheme.
+In the case of minoritarian classes having less than $N$ samples, all the available samples were extracted.
+The training set was generated by extracting patches of $9\times 9$ pixels around the selected points in the training tiles.
+To minimize spatial autocorrelation between the training patches, the corresponding reference data was extracted only for the central $3\times 3$ pixels.
+In this way, we collected a training dataset representative of the entirety of the classification scheme, corresponding to $\sim 1.17\%$ ($1.413.522$ pixels) of the total number of pixels available in the MGRS S2 granule.
+IV-D Design of Experiments
+Validation has been conducted on both datasets considering different multigranularity approaches.
+The reported results include the values of Overall Accuracy (OA) and mean F1 Score (mF1) for both tasks, while for the multispectral dataset also Top-3 and Top-5 metrics are reported.
+We considered different multigranularity modules as comparison, all built on the same backbone architecture for fairness on all the datasets.
+We considered the following approaches:
+  1. 1.
+    Baseline: the baseline with only the fine-grained classification through standard CCE minimization. Selected to evaluate the improvement gained by incorporating the hierarchy.
+  2. 2.
+    HAF [15]: selected as a regularization approach adopting soft labels and geometric constraints that use auxiliary classifiers at each level to ensure consistent predictions with the hierarchy, specifically aiming to reduce the severity of mistakes.
+  3. 3.
+    HRN [8]: included as state-of-the-art benchmark that facilitates hierarchical feature interaction by forcing finer classes to inherit attributes from coarser classes through residual skip connections.
+  4. 4.
+    MGM [31]: included to assess the effectiveness of using knowledge-guided relation graphs and bi-directional interaction structures that enhance coarse-grained features through gradient truncations.
+  5. 5.
+    HiCervix [4]: chosen as benchmark for methods that model semantic correlations within an explicit tree structure using soft labels and recursive probability summation across hierarchical paths.
+  6. 6.
+    ALGS [7]: selected to evaluate the impact of a global semantic distance-based penalty term for adjusting class predictions, ensuring hierarchical consistency when classification is enforced at every level.
+Specifically designed for RS tasks.
+  7. 7.
+    HCL [9]: selected as comparison for representation-based learning that disentangles class-specific embeddings at multiple levels to align them with their semnantic relationships.
+Specifically designed for RS tasks.
+For the VHR task, we considered a ResNet50 backbone pretrained with ImageNet weights [22].
+This choice is due to its deeply stacked convolutional layers and residual connections, making it well-suited for extracting highly localized, fine-grained spatial and morphological features necessary for VHR analysis.
+For the MS Satellite Image Time Series (SITS) classification task, we adopted a different backbone architecture designed to process multispectral and temporal information.
+This backbone consist in an adaptation to the RS scenario of a Swin Transformer [16] and is trained from scratch.
+For a detailed description of the layer topology, please refer to the cited paper [16].
+The choice of this backbone is due to its shifted-window self-attention mechanism that efficiently captures both local details and long-range global contextual dependencies, making it ideal for processing complex multi-scale structural variations.
 We adopted the AdamW optimizer for all the architectures and configurations.
-The metaparameters of each different configuration have been validated considering a Tree-structured Parzen Estimator (TPE) algorithm to efficiently explore the search space.
+The metaparameters of each different configuration have been validated considering a Tree-structured Parzen Estimator (TPE) algorithm to efficiently explore the metaparameters space.
 The learning rate and the weight decay were sampled from log-uniform distributions $U_{\log}([10^{-6},10^{-3}])$.
-Considering the proposed SAHC approach, the weight $\lambda_{HC}$ associated to the HC loss has been sampled from a log-uniform distribution $U_{\log}([10^{-2},10^{2}])$, while the weights $\lambda^{h}$ have been set empirically to $0.3$, $0.2$, and $0.5$, to prioritize the fine-grained classification and reduce the influence of the intermediate and coarse-grained levels.
-To adjust the learning rate during training, we considered a cosine decay schedule.
-The weights associated to each model in the best performing epoch in terms of validation OA% and mF1% were used to conduct the inference on the test set, whose performance metrics are reported in the following section.
+We considered 50 training epochs on both datasets, together with Stochastic Weight Averaging (SWA) to enhance model generalization [19].
+Considering the proposed Semantics-Aware Hierarchical Consensus (SAHC) approach, the weight $\lambda_{HC}$ associated to the HC loss has been sampled from a log-uniform distribution $U_{\log}([10^{-2},10^{2}])$, while the weights $\lambda^{h}$ have been set to $0.5$, $0.2$, and $0.3$ for the VHR dataset, and to $0.4$, $0.2$, $0.1$, and $0.3$ for the MS dataset, to prioritize the fine-grained classification and reduce the influence of the intermediate, coarse-grained, and macro levels.
+We implemented a dynamic consistency warmup schedule for the HC loss weight, $\lambda_{HC}$.
+Specifically, during the training phase, $\lambda_{HC}$ was initially fixed to 0 for the first 5 epochs to allow for stable feature extraction.
+Subsequently, the weight value was linearly ramped up from epoch 5 to 15 until reaching its sampled maximum value, where it remained constant for the duration of the training.
 V Experimental Results
-The experiments focus on validating the proposed SAHC approach over different backbones tasks, and datasets, comparing it to different standard and hierarchical approaches.
+TABLE I: Comparison of the Accuracies obtained by different hierarchical methodologies on the NWPU-RESISC45 dataset. The best results are reported in bold, second best are underlined
+Metric Hierarchy Baseline [22] HRN [8] HAF [15] ALGS [7] HCL [9] MGM [31] HiCervix [4] SAHC
+OA (%) Fine-grained 96.08 96.48 96.16 96.32 95.87 96.11 96.10 96.63
+Intermediate 97.38 97.30 97.35 97.16 97.24 97.19 97.56
+Coarse-grained 97.86 97.86 97.78 97.73 97.84 97.70 97.98
+mF1 (%) Fine-grained 96.08 96.47 96.15 96.33 95.85 96.10 96.09 96.63
+Intermediate 96.74 96.37 96.65 96.26 96.46 96.37 96.89
+Coarse-grained 97.56 97.52 97.47 97.36 97.41 97.37 97.67
+Params (M) 25.8 35.5 25.9 25.9 27.0 35.5 25.9 25.9
+FLOP (T) 6.22 7.03 6.22 6.22 6.51 6.88 6.22 6.22
+The experiments focus on validating the proposed SAHC approach over different datasets, comparing it to other state-of-the-art hierarchical approaches on different levels of the hierarchy.
 We first focus on the results on the specific tasks, i.e., scene classification on the VHR dataset and semantic segmentation of MS SITS.
-Then, we demonstrate the effectiveness of each component of the SAHC approach in an ablation study considering the ELU dataset.
-Finally, we analyze the trained hierarchy projection matrices to identify the key hierarchical relationships learnt by the proposed approach.
+Then, we analyze the importance of the hierarchical weights selection $\lambda^{h}$ on the NWPU dataset.
+Finally, we discuss a visualization of the trained hierarchy projection matrices to identify the key hierarchical relationships learnt by the proposed approach.
 V-A Case I: VHR Scene Classification
-TABLE I: Comparison of the Overall Accuracy obtained by different hierarchical methodologies with different backbone architectures on the NWPU-RESISC45 dataset.
-The best results are reported in bold, second best are underlined.
-The related standard deviation is reported in brackets.
-NWPU-RESISC45
-Backbone Methodology OA (%)
-LWGANet L2 [35] Baseline 92.50 (0.31)
-HAF [18] 93.05 (0.29)
-HRN [20] 93.01 (0.11)
-SAHC (proposed) 94.69 (0.18)
-MobileNet V2 2.5× [34] Baseline 93.70 (0.30)
-HAF 93.67 (0.28)
-HRN 93.58 (0.16)
-SAHC (proposed) 94.19 (0.33)
-EfficientformerV2 S2 [33] Baseline 90.41 (0.38)
-HAF 90.76 (0.46)
-HRN 91.16 (0.44)
-SAHC (proposed) 91.72 (0.14)
-Table I shows the OAs of the LWGANet, MobileNet, and EfficientFormer on the NWPU-RESISC45 dataset.
-The proposed SAHC achieves the best results, showing a higher OA% with all the different backbones.
-The HAF and HRN hierarchical modules show improvements only in the case of the transformer-based backbone (i.e., EfficientFormer) and the hybrid backbone (i.e., LWGANet), while the baseline approach without hierarchical modules performs better with the CNN-based approach (i.e., MobileNet).
-In contrast, the proposed SAHC shows an increment in OA% of $\sim 1.2\%$ compared to the baseline approach with both the transformer-based and the CNN-based architectures, while exhibits a greater improvement when considering the hybrid backbone, with an additional $\sim 2.2\%$ increase in mF1%.
-These findings highlight the contribution of the hierarchical module.
-Indeed, its application to the hybrid architecture results in higher performance compared to its application to the CNN model.
-This result is particularly interesting given that the baseline CNN model outperforms the LWGANet baseline.
-It is also worth noting that, in most cases, the SAHC shows a smaller variability compared to its competitors, ensuring stabler results across different random initializations of the architecture weights.
+The performance of the proposed SAHC method on the NWPU-RESISC45 dataset is summarized in Table I, providing a comparative analysis against state-of-the-art methodologies and the non-hierarchical baseline.
+Given the balanced class distribution characteristic of the NWPU-RESISC45 dataset, the OA and mF1 are aligned across all evaluated models.
+From the table, one can see that the integration of hierarchical information consistently improves classification performance over the baseline.
+Indeed, the majority of the evaluated hierarchical architectures outperform their non-hierarchical counterpart (with the exception of Hierarchical Contrastive Learning (HCL)), confirming that the exploitation of semantic relationships allows for a more robust distinction between fine-grained RS classes.
+Among the compared methodologies, the proposed SAHC achieves the highest performances at the fine-grained level, with an OA of 96.6396.63% and an mF1 of 96.6396.63%.
+Notably, the good performance of SAHC extends across the entire taxonomic hierarchy.
+At the intermediate and coarse-grained levels, the proposed method achieves OAs of 97.5697.56% and 97.9897.98%, respectively, consistently outperforming the other frameworks.
+The second-best overall performance is obtained by the HRN method, which reaches a fine-grained OA of 96.4896.48%.
+However, this comparison also highlights the computational overhead introduced by the different architectures.
+The HRN method achieves its high accuracy at the cost of architectural efficiency.
+Indeed, the introduction of the Granularity-Specific Block (GSB) leads to a significant increase in parameters and computational overload.
+In contrast, the proposed SAHC maintains the same computational footprint as the baseline architecture.
+This efficiency is shared by other methods such as Adaptive Label Granularity Selection (ALGS), HAF, and HiCervix, where the bottleneck remains the backbone itself.
+In the specific case of SAHC, the architectural overhead is due to the three auxiliary heads and their associated projection matrices, both of which are negligible relative to the dimension of the backbone.
+When focusing strictly on architectures that do not introduce additional complexity, the ALGS method provides the highest performance after SAHC, though it remains lower in terms of OA.
 V-B Case II: MS SITS Semantic Segmentation
-TABLE II: Comparisons of the accuracies obtained by different hierarchical methodologies with different backbone architectures on the ELU and HRLC-CCI datasets.
-The best results are reported in bold, second best are underlined.
-The related standard deviation is reported in brackets.
-ELU Dataset HRLC-CCI Dataset
-Backbone
-Methodology
-mF1 (%)
-OA (%)
-mF1 (%)
-OA (%)
-Swin Transformer [37] Baseline (no hierarchy) 33.42 (0.44) 68.10 (0.65) 64.03 (1.06) 83.30 (1.22)
-HAF [18] 32.97 (0.57) 68.15 (0.56) 64.26 (0.41) 83.35 (0.68)
-HRN [20] 30.48 (0.86) 67.72 (1.98) 63.80 (0.38) 83.05 (0.49)
-SAHC (proposed) 34.57 (0.12) 71.25 (0.25) 66.49 (0.34) 84.75 (0.47)
-Transformer [36] Baseline (no hierarchy) 18.77 (0.29) 54.79 (0.87) 61.14 (0.62) 81.74 (0.72)
-HAF 19.04 (0.28) 55.69 (0.88) 61.12 (0.81) 81.37 (1.03)
-HRN 20.97 (0.58) 57.48 (2.31) 60.46 (1.05) 81.80 (0.60)
-SAHC (proposed) 25.63 (0.41) 65.94 (0.74) 64.07 (0.45) 83.30 (0.44)
-ConvLSTM [38] Baseline (no hierarchy) 25.82 (0.95) 62.59 (0.70) 65.19 (0.77) 84.32 (0.33)
-HAF 26.11 (0.92) 62.40 (0.57) 65.64 (0.84) 84.42 (0.68)
-HRN 25.53 (1.03) 62.36 (1.47) 65.18 (0.31) 84.19 (0.58)
-SAHC (proposed) 26.48 (0.35) 66.74 (0.42) 65.38 (0.40) 85.08 (0.28)
-The second study case analyzes the semantic segmentation task, and is characterized by two different datasets with varying hierarchical complexity and differing label quality.
-The ELU dataset employs weak labels due to its focus on land use rather than land cover.
-Conversely, the HRLC-CCI dataset is designed for pixel-based LC classification, providing more precise and location-specific labels.
-Table II shows the performance of the three considered backbones on both the ELU and the HRLC-CCI dataset in terms of mF1% and OA%.
-The proposed approach outperforms the other methods on the ELU dataset, which exhibits stronger hierarchy and weaker labels than the HRLC-CCI dataset.
-Notably, the pixel-based transformer without contextual information shows an improvement of about $7\%$ in terms of mF1% compared to the non-hierarchical baseline, while it also outperforms the other hierarchical approaches.
-In the case of the hybrid Swin Transformer and the ConvLSTM, the performance gap tends to decrease, while the adoption of the hierarchy and the consensus approach still shows an improvement.
-This result underlines the robustness of the proposed method to noisy labels, as the trained hierarchical matrices can learn soft hierarchical relationships directly from the data.
-In contrast, the compared methodologies rely on a hard hierarchy definitions, which might lead to suboptimal results in the case of labels that do not accurately represent the observed samples.
-The same behavior can be observed when analyzing the results on the HRLC-CCI dataset, where the hierarchy becomes weaker, but the real targets are better represented by the labels at the pixel level.
-In this case, the proposed approach performs similarly to the other methodologies, as the hierarchy is less noisy and the mapping more robust.
-Nonetheless, SAHC still outperforms the other methodologies, showing good generalization across the different backbones and scenarios.
-In our experiments, SAHC performed worse than HAF in terms of mF1% score only in the case of the ConvLSTM backbone, although the performance remains comparable.
-However, the proposed approach still shows a higher OA%, with reduced scores variability across different seeds.
+TABLE II: Comparison of the Accuracies obtained by different hierarchical methodologies on the ELU dataset. The best results are reported in bold, second best are underlined
+Metric Hierarchy Baseline [22] HRN [8] HAF [15] ALGS [7] HCL [9] MGM [31] HiCervix [4] SAHC
+OA (%) Fine-grained 36.39 36.41 36.63 35.90 37.26 35.42 36.92 38.08
+Intermediate 46.60 41.58 46.21 48.81 45.99 47.91 49.68
+Coarse-grained 57.13 57.89 56.43 59.58 56.30 58.27 59.90
+Macro 76.34 78.82 75.49 78.36 75.44 77.39 78.35
+Top-3 (%) Fine-grained 56.48 57.57 55.86 54.75 56.94 55.83 56.98 58.96
+Intermediate 71.01 63.61 69.27 71.59 70.13 71.31 73.48
+Coarse-grained 82.44 81.15 80.90 82.76 81.66 82.29 84.31
+Macro 96.77 96.36 97.10 97.30 96.99 97.50 97.23
+Top-5 (%) Fine-grained 64.57 66.61 64.21 63.11 65.54 64.89 65.48 67.84
+Intermediate 80.68 74.55 78.72 80.60 79.92 80.66 82.80
+Coarse-grained 90.67 88.88 90.07 90.69 90.18 90.93 91.98
+Macro -— -—
+mF1 (%) Fine-grained 32.02 32.38 32.12 31.63 33.27 31.60 32.59 34.25
+Intermediate 39.25 27.92 38.60 41.03 39.19 39.86 42.59
+Coarse-grained 52.64 46.42 51.44 54.80 51.98 53.23 55.34
+Macro 66.96 70.64 67.12 69.47 66.48 68.95 69.12
+Params (M) 4.7 16.2 4.8 4.8 5.8 16.2 4.7 4.8
+FLOP (T) 2.22 2.38 2.22 2.22 2.41 2.32 2.22 2.22
+To further evaluate the robustness of the proposed framework, an extensive analysis was conducted on the highly complex ELU dataset.
+Unlike the previous scenario, this dataset consists of lower-resolution multispectral images (10m10m/pixel) mapped to a complex taxonomy of $90$ classes distributed across four hierarchical levels.
+The results of this analysis are reported in Table II.
+The complexity of the ELU dataset, combined with the potential for semantic noise or inconsistencies in the CLC LU, makes this scenario particularly challenging.
+Standard hierarchical models that incorporate a rigid taxonomy might struggle to adapt to these inconsistencies.
+The proposed SAHC mitigates this limitation by leveraging a data-driven taxonomy, allowing the architecture to dynamically adapt even when the predefined hierarchy is not perfectly aligned.
+This advantage can be observed in the OA and mF1 shown in the Table.
+Indeed, SAHC achieves the best performance at the fine-grained level with an OA of 38.0838.08% and an mF1 of 34.2534.25%.
+These results are also confirmed at the intermediate (OA 49.6849.68%, mF1 42.5942.59%) and coarse-grained levels (OA 59.9059.90%, mF1 55.3455.34%), consistently better performing with respect to other state-of-the-art methodologies like HiCervix and HCL.
+At the macro level, SAHC demonstrates competitive performance, ranking after only to the HCL and HAF methods by a marginal gap.
+To further evaluate the ability of the models to capture the semantic relationships, we analyzed the Top-3 and Top-5 accuracy metrics across all the levels of the hierarchy (see Table II).
+These metrics indicate whether the ground-truth class is among the highest-confidence predictions of the model, measuring the capacity of the network to maintain semantic proximity even when the absolute Top-1 prediction is incorrect.
+The proposed SAHC approach demonstrates the highest performance across almost all granularity levels.
+Specifically, SAHC achieves the highest Top-3 accuracies for the fine-grained, intermediate, and coarse-grained levels, reaching a 58.9658.96%, 73.4873.48%, and 84.3184.31%, respectively.
+This represents a large improvement against the HRN, which ranked second in the VHR case, showing a margin of 1.391.39% at fine-grained level, while also outperforming the RS-oriented competitor HCL by 1.891.89% and 1.551.55% in the intermediate and coarse levels, respectively.
+At the macro level, Top-3 performance saturates for all evaluated models.
+In this aggregated space, SAHC remains robust (97.2397.23%), performing comparably to HCL (97.3097.30%) and HiCervix (97.5097.50%).
+Because the total number of target classes at the macro level is small, Top-3 classification becomes a less discriminative metric.
+For the Top-5 accuracy, SAHC maintains the overall best performance, outperforming the baseline and the compared methodologies across the evaluated granularities.
+At the fine-grained level, SAHC achieves 67.8467.84%, surpassing the HRN by 1.231.23%.
+This performance gap widens at the intermediate level, where SAHC reaches 82.8082.80% compared to HRN 80.6880.68%.
+At the coarse-grained level, SAHC achieves 91.9891.98%, outperforming the HiCervix by 1.051.05%.
+Note that Top-5 metrics are omitted at the macro level, as the limited number of classes at this highest taxonomy level renders the Top-5 metric uninformative.
+The improvement observed in the Top-3 and Top-5 accuracies can be attributed to the design of the SAHC architecture.
+When the network struggles to isolate the exact fine-grained class, the hierarchical consensus mechanism, which ensures probabilistic alignment across different hierarchical levels, guarantees that the subsequent highest-confidence predictions belong to semantically related classes within the same hierarchical branch.
+This minimizes the divergence between the projected predictions and the overall committee consensus, reducing severe misclassifications.
+Figure 5: Qualitative example of the predictions at the fine-grained level on the ELU dataset obtained by the considered hierarchical methodologies using the ResNet50 backbone: (a) true color (RGB) monthly composite for August 2020, (b) CLC reference map, (c) non-hierarchical baseline (d) HRN method, (e) HAF method, (f) ALGS method, (g) HCL method, (h) MGM method, (i) HiCervix method, and (j) the proposed SAHC method.
+Finally, consistent with the previous analysis, SAHC achieves these state-of-the-art results without sacrificing computational efficiency.
+While architectures like HRN and multi-granularity module (MGM) impact on the computational efficiency, SAHC keeps the computational footprint nearly identical to the baseline model.
 Lastly, Fig.
-4 shows a visual comparison of the different hierarchical methodologies considering the HRLC-CCI dataset and the Swin Transformer backbone.
-From the qualitative results, one can see that the SAHC approach better represents the “Tree” classes, also showing a lower pixel-level noise.
-Figure 4: Example of qualitative results at the fine-grained level on the HRLC-CCI dataset obtained by the considered methods using the Swin Transformer backbone.
-V-C Ablation Study
-TABLE III: Ablation study on the Swin backbone with the ELU dataset.
-The classification metrics are reported with the related standard deviation in brackets.
-The best results are reported in bold, second best are underlined.
-Multi-level
-Classifiers Hierarchical
-Consensus
-(Training) Hierarchical
-Consensus
-(Inference) mF1 (%)
-OA (%)
-33.42 (0.44) 68.10 (0.65)
-33.48 (0.59) 68.10 (1.06)
-34.10 (0.41) 67.83 (1.00)
-33.48 (0.37) 68.39 (0.33)
-34.57 (0.12) 71.25 (0.25)
-The performance of each component in the proposed hierarchical approach is further analyzed in the ablation study presented in Table V-C.
-In this study, we evaluate the Swin backbone across five distinct configurations on the ELU dataset:
+5 provides a visual comparison of the evaluated hierarchical methodologies on a small portion of the ELU dataset.
+From the qualitative results, one can see that the SAHC approach retrieves the most spatially-consistent land-cover maps, characterized by a significant reduction in classification artifacts compared to the competing approaches.
+This is particularly evident in the southern sector of the study area.
+Indeed, while in this specific case most networks misclassify “transitional woodland/shrub” as “inactive quarries” due to sparse vegetation signatures, SAHC correctly identifies the vegetative cover, showing robustness in classification.
+These advantages are further evident in Fig.
+6.
+From a broader perspective, scaling a RS classification problem to an extensive taxonomy of $90$ classes inherently introduces severe spectral confusion and extreme class imbalance.
+In this context, the observed improvements in the mF1 score are significant.
+Indeed, they demonstrate that the proposed SAHC mitigates the ambiguities typical of such label space, maintaining robust and granular predictions where traditional approaches fail.
+Figure 6: Qualitative example of the predictions at the fine-grained level on the ELU dataset obtained by the considered hierarchical methodologies using the ResNet50 backbone: (a) true color (RGB) monthly composite for August 2020, (b) CLC reference map, (c) non-hierarchical baseline (d) HRN method, (e) HAF method, (f) ALGS method, (g) HCL method, (h) MGM method, (i) HiCervix method, and (j) the proposed SAHC method.
+V-C Sensitivity Analysis
+Figure 7: Sensitivity analysis of the hierarchical weights $\lambda^{h}$ on the VHR dataset.
+The performance surface is visualized as a probability simplex, where each point represents a weight combination for the three hierarchical levels sampled at a step size of $0.1$.
+The simplex heatmap shows the optimal configuration as $\lambda^{1}=0.5$, $\lambda^{2}=0.2$, $\lambda^{3}=0.3$.
+To evaluate the robustness of the proposed hierarchical architecture, we conducted a sensitivity analysis on the VHR dataset to assess how model performance varies with the selection of different granularity weights $\lambda^{h}$.
+The search space was constrained such that the weights sum to one ($\sum_{h=1}^{H}\lambda^{h}=1$), with each $\lambda^{h}$ sampled using a step size of $0.1$.
+Since the architecture adopts three hierarchical levels, the resulting parameter space can be visualized as a probability simplex that illustrates the various weight combinations across the different granularities.
+Fig.
+7 illustrates a heatmap reporting the performance across all evaluated configurations of the three layer weights.
+As demonstrated in the figure, the optimal combination is achieved at $\lambda^{1}=0.5$, $\lambda^{2}=0.2$, $\lambda^{3}=0.3$.
+This optimal configuration justifies the hypothesis that priority should be allocated to both the fine-grained classification task and the coarsest class.
+Nevertheless, the analysis also reveals the presence of several local maxima across the performance surface, which are generally skewed toward configurations in favor of the finest-level weight.
+V-D Ablation Study
+TABLE III: Ablation study on the ResNet50 backbone with the
+NWPU-RESISC45 dataset. The best results are
+in bold, second best are underlined
+Multi-Heads $L_{HC}$ $\hat{\mathbf{p}}_{HC}$ $L^{HC}_{CCE}$ Hierarchy OA (%) mF1 (%)
+Fine 96.08 96.08
+Inter.
+Coarse
+Fine 96.44 96.43
+Inter. 97.41 96.68
+Coarse 97.92 97.58
+Fine 96.49 96.48
+Inter. 97.44 96.68
+Coarse 97.94 97.57
+Fine 96.52 96.53
+Inter. 97.41 96.64
+Coarse 97.76 97.33
+Fine 96.49 96.51
+Inter. 97.48 96.71
+Coarse 97.90 97.48
+Fine 96.63 96.63
+Inter. 97.54 96.86
+Coarse 97.97 97.66
+The performance contribution of each component in the proposed hierarchical approach is analyzed in the ablation study presented in Table III.
+In this study, we evaluate the ResNet50 backbone across six different hierarchical configurations on the NWPU-RESISC45 dataset:
   1. 1.
-    Baseline: the baseline approach with no multi-level classifiers;
+    Baseline: the standard approach with no multi-level classifiers;
   2. 2.
-    Multi-level: the baseline approach augmented with additional classification heads for the coarse-grained and intermediate classes;
+    Multi-level: the baseline approach modified with additional classification heads for the coarse-grained and intermediate classes.
   3. 3.
-    Multi-level with HC inference: the multi-level approach with consensus applied solely during inference (considering user-defined hierarchies);
+    Multi-level with HC inference $\hat{\mathbf{p}}_{HC}$: the multi-level approach with consensus applied solely during inference (considering user-defined hierarchies).
   4. 4.
-    Multi-level with $\mathcal{L}_{HC}$ training: the multi-level approach trained with the self-supervised consensus loss, excluding consensus-based classification during inference;
+    Multi-level with $\mathcal{L}_{HC}$ training: the multi-level approach trained with the self-consistent consensus loss, excluding consensus-based classification during inference.
   5. 5.
-    SAHC: the proposed approach, which incorporates both the self-supervised consensus loss and consensus inference at multiple levels.
-Figure 5: Heatmap of the estimated hierarchy log-joint matrices considering the Swin Transformer backbone on the ELU dataset from fine-grained classes to intermediate classes.
-The blue and green rectangles identify the expected hierarchical aggregation of the fine-grained classes at the intermediate and coarse levels, respectively.
-The table results indicate that incorporating multi-level classifiers into the baseline does not improve the overall performance.
-In fact, it leads to increased variability in the results, showing a higher variance and making the outcomes more unstable.
-However, the introduction of the HC inference improves the representation of fine-grained classes in terms of mF1%, also when considering the user-defined hierarchy in the multi-level approach with HC inference.
-On the contrary, the multi-level approach with $\mathcal{L}_{HC}$ training does not improve the performance, as the JSD acts on the consensus across the layers and not on the fine-grained classification itself.
-This is instead addressed by the proposed SAHC, which combines the self-supervised consensus loss with consensus inference at multiple hierarchy levels.
-By allowing the trained matrices to learn the relationships from the data, the hierarchical consensus exhibits a sharp improvement in fine-grained classification, outperforming all other configurations.
-It is also important to note that the introduction of the JSD and the hierarchical consensus in inference leads to lower variance values, showing a more consistent fine-grained classification across different weights initializations compared to the other approaches.
-V-D Hierarchy Projection Matrices Analysis
-Figure 6: Heatmap of the estimated hierarchy log-joint matrices considering the Swin Transformer backbone on the HRLC-CCI dataset from fine-grained classes to intermediate classes.
-The blue and green rectangles identify the expected hierarchical aggregation of the fine-grained classes at the intermediate and coarse levels, respectively.
-To better visualize the impact of the hierarchy projection matrices on classification results, we analyze the estimated hierarchy projection matrices of the Swin Transformer on the ELU and HRLC-CCI datasets.
-The visualization on the NWPU dataset is not reported as the hierarchical mappings show a trivial result, with perfect mapping of the defined hierarchical levels (mostly due to the simpler classification task).
-Figure 5 presents a heatmap of the log-joint matrix between intermediate and fine-grained levels, highlighting the hierarchical structure captured by the proposed methodology on the ELU dataset.
-Note that the hierarchy projection matrix was uniformly initialized before training, with no prior knowledge of the user-defined hierarchical structure of the problem.
-The results show that the proposed approach effectively learns the hierarchy in the dataset.
-Specifically, the blue rectangles in the heatmap correspond to the user-defined expected mappings, from fine-grained classes to intermediate classes.
-Here, the weights associated to these mappings are highlighted, confirming that the data-driven estimated mappings are correctly following the user-defined hierarchies.
-Then, green rectangles indicate distant mappings to coarse-grained classes, revealing a similar pattern even though the matrix focuses on mappings between fine-grained and intermediate levels.
-It is important to note that separate hierarchy joint matrices are trained for the different hierarchical mappings in the proposed approach.
-Similarly, Figure 6 depicts the heatmap associated with the HRLC-CCI dataset.
-As observed in the ELU case, the proposed methodology correctly extracts the hierarchy within the problem analyzed, emphasized by the blue rectangle showing the intermediate aggregation.
-However, the mappings appear more dispersed compared to the more complex ELU scenario, mostly due to the simpler hierarchical constraint leading to a smaller impact.
-Nonetheless, the trainable log-joint matrices allow soft-hierarchical mappings that were not considered in the initial user-defined hierarchy.
-One example can be observed in the weights associated to “Grassland vegetation aquatic or regularly flooded areas”, which shows an active semantic relationship with the “Open water” class, which is not mapped by the imposed user-defined hierarchy.
-This mapping highlights the importance of a data-driven approach, in contrast with the superimposition of a rigid hierarchical structures.
+    Multi-level with $\mathcal{L}_{HC}$ training and HC inference $\hat{\mathbf{p}}_{HC}$: the multi-level approach combining both the self-supervised consensus training and consensus-based classification during inference.
+  6. 6.
+    SAHC: the proposed full approach, which incorporates the self-supervised consensus loss, consensus inference, and the supervised consensus loss ($\mathcal{L}^{HC}_{CCE}$) at multiple levels.
+As observed in Table III, the integration of the proposed modules consistently enhances the performance of the network.
+The transition from the baseline to the simple multi-level architecture shows an increase of the fine-grained OA from 96.0896.08% to 96.4496.44%.
+This demonstrates that simply forcing the network to learn coarse and intermediate weighted representations provides regularization for the fine-grained task.
+Introducing the HC inference constraint or the self-consistent loss individually provides further performance gains.
+Notably, using only the self-consistent consensus training leads to good fine-grained OA of 96.5296.52% and an mF1 of 96.5396.53%, suggesting that enforcing hierarchical consistency during the learning phase is effective.
+Interestingly, combining both modules creates a trade-off, reducing marginally the fine-grained accuracy while improving the intermediate OA to 97.4897.48%.
+Finally, the introduction of the supervised consensus loss in the complete SAHC framework provides the best trade-off between training and inference constraints.
+The full SAHC configuration achieves the highest metrics across all hierarchical levels, achieving the best fine-grained OA of 96.6396.63%.
+This confirms that the cooperation between self-consistent alignment, structured inference, and supervised hierarchical correction maximises the discrimination capabilities of the network.
+V-E Hierarchy Projection Matrices Analysis
+Figure 8: Heatmap of the estimated joint probability matrix $\mathbf{J}^{h_{2},h_{1}}$ considering the ResNet50 backbone on the ELU dataset from fine-grained classes to intermediate classes.
+The cyan rectangles identify the expected hierarchical aggregation of the fine-grained classes at the intermediate level.
+To better visualize the impact of the hierarchical projection matrices on classification results, we analyze the estimated joint probability matrix $\mathbf{J}^{h_{2},h_{1}}$ representing the learned hierarchy between the fine-grained and intermediate levels for the ELU dataset.
+The corresponding visualization on the NWPU dataset is omitted, as its simpler classification task leads to an expected result where the learned mappings perfectly align with the predefined hierarchical levels.
+Figure 8 presents a heatmap of this joint probability matrix, illustrating the hierarchical structure captured by the proposed SAHC methodology on the ELU dataset.
+The visualization confirms that the proposed approach effectively learns the hierarchy in the dataset, correctly identifying the semantic aggregations between the levels.
+Specifically, the cyan rectangles in the heatmap correspond to the user-defined expected mappings, from fine-grained classes to intermediate classes.
+The higher weights within these regions confirm that the data-driven estimated mappings successfully respect the intended user-defined taxonomy.
+In addition to these rigid hierarchies, the network also captures important semantic relationships between different coarser groups.
+Indeed, the blue rectangles indicate cross-groups relationships related to to coarser-grained classes.
+Because separate joint matrices are trained for different hierarchical transitions in the proposed architecture, the model is able to discern soft-hierarchical mappings that were strictly excluded from the initial user-defined hierarchy.
+For instance, the learned weights reveal an active semantic relationship between “water basins” and “aquaculture in the marine environment”, as well as between “continuous urban fabric” and “road networks”.
+These mappings highlight the importance of a data-driven approach.
+Indeed, rather than superimposing a rigid and potentially flawed hierarchical structure, the model dynamically adapts to the true semantic distributions present in the RS image.
 VI Conclusion
-In scene classification and semantic segmentation, the hierarchical relationship among the target classes is often under-exploited, resulting in the exclusion of valuable semantic information from the decision-making process.
+In scene classification and semantic segmentation RS tasks, the hierarchical relationship among the target classes is often under-exploited, resulting in the exclusion of valuable semantic information from the decision-making process.
 To address this limitation, this paper proposes the SAHC approach that leverages the hierarchical structure inherent in class types to improve the robustness and the accuracy of the classification.
 We introduced hierarchical classification branches, each designed to handle a specific hierarchical classification task, ranging from coarse-grained to fine-grained categories.
-Trainable hierarchy log-joint matrices are then defined to project the predictions across different hierarchical levels, following a self-supervised data-driven approach which may not align with the rigid user-defined mappings between different hierarchy levels.
+Trainable hierarchy log-joint matrices are then defined to project the predictions across different hierarchical levels, following a self-consistent data-driven approach which may not align with the rigid user-defined mappings between different hierarchy levels.
 This allows the implementation of a consensus mechanism, where predictions from multiple hierarchical levels are fused to perform the final decision.
-By progressively guiding the classification process from broader to specific categories, we can optimize the consensus, ensuring alignment among predictions at different hierarchical levels in a self-supervised fashion.
-The results show that data-driven hierarchies improve the performance at the finest level compared to directly using the user-defined hierarchies, contrarily to the performance drop usually observed in the literature when adding additional classification layers for different hierarchical levels after the backbone.
+By progressively guiding the classification process from broader to specific categories, we can optimize the consensus, ensuring alignment among predictions at different hierarchical levels in a self-consistent fashion.
 The proposed methodology is both lightweight and scalable, making it easily adaptable to different architectures and datasets by simply reorganizing their structure in a hierarchical manner.
-We evaluated the performance of the proposed approach on three different RS datasets.
+Extensive experimental analysis demonstrated the effectiveness of the proposed approach on two different RS datasets, representing a VHR and a MS classification tasks.
 Indeed, the proposed approach is particularly relevant in VHR image and RS MS SITS classification, where the semantic structure of the data and available datasets support an explicit hierarchical representation of the task.
-The experimental results reveal that the proposed methodology is able to exploit both complex and simple hierarchies, demonstrating the best performances in terms of OA% and mF1% scores across experiments at the fine-grained hierarchical level.
+The experimental results reveal that the proposed methodology is able to exploit both complex and simpler hierarchies, demonstrating the best performances in terms of OA% and mF1% scores across experiments across all hierarchical levels.
 As expected, the improvements are particularly significant when complex hierarchies can be exploited within the RS data.
-The proposed method is general, showing top performance scores across different tasks and architectures.
+The proposed method is general, showing top performance scores across different tasks.
+Furthermore, the self-consistent consensus loss effectively constrained the hierarchical levels to converge toward a unified prediction, reducing hierarchical violations in the final classification maps.
 A limitation of the proposed method is the reliance on a predefined hierarchy within the labels to train the self-supervised hierarchical consensus.